### v2

Abstract
This paper introduces ExpertLongBench, an expert-level benchmark containing 11 tasks from 9 domains that reflect realistic expert workflows and applications. Beyond question answering, the application-driven tasks in ExpertLongBench demand long-form outputs that can exceed 5,000 tokens and strict adherence to domain-specific requirements. Notably, each task in ExpertLongBench includes a rubric, designed or validated by domain experts, to specify task requirements and guide output evaluation. Furthermore, we propose Clear, an evaluation framework that supports accurate evaluation of long-form model outputs in our benchmark. To achieve fine-grained, expert-aligned evaluation, Clear derives checklists from both model outputs and references by extracting information corresponding to items in the task-specific rubric. Checklist items for model outputs are then compared with corresponding items for reference outputs to assess their correctness, enabling grounded evaluation. We benchmark 11 large language models (LLMs) and analyze components in Clear, showing that (1) existing LLMs, with the top performer achieving only a 26.8% F1 score, require significant improvement for expert-level tasks; (2) models can generate content corresponding to the required aspects, though often not accurately; and (3) accurate checklist extraction and comparison in Clear can be achieved by open-weight models for more scalable and low-cost usage.
footnotetext: Co-first authors. **footnotetext: Correspondance to {jieruan,inair,caoshuy,wangluxy}@umich.edu
1 Introduction
There is growing interest in applying large language models (LLMs) to specialized applications that demand domain-specific expertise, such as student tutoring [1], legal case summarization [2], and medical diagnosis [3]. These expert-level tasks pose greater challenges because they require specialized knowledge and strict adherence to domain-specific standards, which are typically acquired through advanced education and professional experience. Furthermore, many real-world expert-level tasks require understanding lengthy inputs [4] and generating complex and nuanced long-form outputs [5, 6].
Existing expert-level benchmarks such as MMLU [7] and GPQA [8] prioritize ease of evaluation of multiple-choice or short-form answers at the expense of alignment with realistic expert applications. Though ExpertQA [9] assesses longer-form, domain-specific outputs, it remains centered on QA scenarios rather than end-to-end expert workflows. In addition to task scope, a key limitation is the lack of evaluation methods tailored to the specific requirements of each task. WildBench [10] creates checklists to evaluate model responses to user queries, yet their model-generated checklists fall short in capturing domain-specific requirements, which often demand expert knowledge to construct. Finally, a broader challenge is the lack of reference outputs in expert-level benchmarks with long-form responses, leading to ungrounded evaluations and oversight of recall estimation.
To address these gaps, this paper first introduces ExpertLongBench, a multi-disciplinary, expert-level benchmark with tasks that require long-form outputs. ExpertLongBench comprises 11 tasks across 9 domains, including law, material science, education, healthcare, chemistry, biology, medicine, finance, and cybersecurity, with 1050 samples in total (Tab. 1). Our benchmark enables more comprehensive assessment of a model’s problem-solving capabilities by including domain-specific end-to-end applications, such as drafting legal briefs or clinical notes, with maximum input and output lengths exceeding 200K and 5K tokens, significantly longer than existing datasets (see comparison in Appendix LABEL:ap:comparison). The completion of these tasks is time-consuming, even for human experts. For example, it takes over 10 hours for a proficient legal practitioner to summarize a complex legal case, as they need to follow the docket and review tens to hundreds of court filings [5]. Moreover, ExpertLongBench provides expert-written reference outputs for each task, facilitating grounded evaluation.
Figure 1: Pipeline of Clear. The example shown is from task T1: multi-document legal case summarization. The checklist mapper takes as input the model output (or human-written reference) and extracts checklist items according to the rubric. Checklists of the model output and the reference are compared at the item level, and the results are subsequently aggregated into the final scores of precision, recall, F1, and accuracy.
To ensure more accurate evaluation of model performance on expert-level tasks, we further propose ClearCheckList-based Expert-level Assessment with Rubric (Fig. 1). We collaborate with domain experts to develop and validate a detailed evaluation rubric for each task. Unlike subjective criteria, such as “usefulness” [9] and “helpful information” [10], our rubric specifies the essential elements required in the output for each application, enabling fine-grained evaluation that closely align with domain-specific requirements. For example, a well-written summary of legal case documents intended for lawyers must accurately identify the cause of action, the relevant statutory or constitutional basis, and the remedy sought. Based on the rubric, Clear breaks down model responses and reference outputs into itemized checklists. For each item, our framework compares the corresponding information extracted from the model’s output against that of the reference output, thereby ensuring a grounded and objective checklist-based assessment. [^1] The benchmark leaderboard and links to code and public data: https://huggingface.co/spaces/launch/ExpertLongBench.
We evaluate the performance of 11 frontier LLMs on ExpertLongBench, including both open-weight and proprietary model families. Our findings are: $\bullet$ The best model achieves an average F1 score of 26.8% across all tasks, revealing that the end-to-end expert-level tasks in ExpertLongBench present significant challenges for existing LLMs. $\bullet$ Despite a low overall output quality, models can generate content that matches over 67% of the aspects required in the checklist, suggesting a risk of models producing content that appears expert-aligned but is potentially misleading due to incorrect generation. $\bullet$ Clear can be configured with open-weight models for cost-effective and scalable benchmarking, as evidenced by a Pearson correlation of 0.88 between the scores assigned by Qwen2.5-72B and GPT-4o for checklist-based evaluation.
2 Related Work
Evaluating LLMs on Expert-Level Knowledge.
Previous studies have assessed domain-specific knowledge using multiple-choice questions (e.g., MMLU [7], GPQA [8]) or short-answer questions (e.g., AGIEval [11], SciEval [12], HLE [13], OlympiadBench [14]). Their scope remains constrained to exam-style questions, failing to reflect domain-specific tasks in real world that typically require long-form generation.
ExpertQA curates open-ended questions written by domain experts [9]. However, the answers to these questions are typically around 100 words, as the questions are formulated to address experts’ specific knowledge queries rather than addressing the full scope of an end-to-end workflow. Additionally, the evaluation method relies on human judgments of answer usefulness, which are inherently subjective, coarse-grained, and lack scalability.
ExpertLongBench addresses these gaps by introducing diverse, expert-level tasks that align with complete workflows; incorporating tasks requiring long-form generation and featuring lengthy inputs that are typical in practical use cases; and providing rubric-based evaluation for fine-grained, domain-specific assessment.
Table 1: Benchmark statistics. : rubric is developed by experts; otherwise, it is created by refining and expanding upon established evaluation protocols; : task data is held privately. For each task, we report whether the task data is newly created (✓) or adapted from previous work; the average number of checklist items in each sample (#Rubric); and the average length of the input (#Input) and human reference (#Reference). Several of these tasks feature significantly longer inputs and references compared to existing domain-specific datasets (see Appendix LABEL:ap:comparison).
Task Description New Sample num #Rubric #Input #Reference
T1LegalMDS * Multi-Document case Summarization 100 26 113,745 1,778
T2LegalSFG * Statement of Fact Generation 100 41 187,302 5,155
T3MaterialSEG * Synthesis Explanation Generation 50 6 199 125
T4EduPAE Pedagogical Alignment Evaluation 100 7 277 60
T5EduFG *† Feedback Generation 100 19 666 139
T6HealthCNG Clinical Note Generation 100 29 1,304 479
T7ChemMDG * Molecule Description Generation 100 6 112 197
T8BioPDG * Protein Description Generation 100 5 111 274
T9MedicalDR Diagnosis & Reasoning 100 11 1,335 429
T10FinanceESG ESG report generation 100 27 69,297 280
T11CyberRDG Risk Description Generation 100 6 555 157
Evaluating Long-Form Generations.
Beyond traditional shallow n-gram overlap-based metrics [15, 16], BLEURT [17] and BERTScore [18] use contextualized embeddings to assess semantic similarity between generated and reference texts. For more granular evaluation, recent research explores fact decomposition-based methods, extracting atomic facts from generated and reference texts for comparison using NLI models or LLMs. [19, 20]. Nevertheless, the fact decomposition process lacks task-specific considerations for fact granularity, which can result in inconsistent evaluations [21].
Another line of work, commonly termed “LLM-as-a-judge”, prompts LLMs to assess generated outputs against general, high-level criteria such as fluency, coherence, and relevance [22]. To enhance task specificity, checklist-based evaluations are incorporated into WildBench [10] and BiGGenBench [23], which evaluate outputs against instance-specific checklists derived from task requirements. However, the absence of grounding for these checklists in reference outputs can introduce coarse and vague criteria, often leading to subjective assessments and hindering the accurate evaluation of content coverage and relevance.
Our evaluation framework, Clear, builds on these directions by extending fact decomposition and checklist-based evaluation. We derive task-specific checklists from an expert-designed rubric, thereby adapting prior checklist-based methods to better support expert-level tasks. For each checklist item, we extract relevant facts from the model outputs and/or references and compare them to compute item-level scores, enabling fine-grained and objective evaluations.
3 Benchmark Construction with Multi-disciplinary Expert-Level Tasks
3.1 Overview
We introduce ExpertLongBench, a multi-disciplinary benchmark designed to evaluate the capabilities of LLMs on real-world expert-level tasks that require long-form inputs and outputs. ExpertLongBench features a fine-grained evaluation rubric tailored to reflect the rigorous standards and nuanced requirements of expert domains. The benchmark covers 11 expert-level tasks as shown in Tab. 1, including six tasks with newly collected data featuring unique challenges such as super-long input, where experts typically spend over 10 hours to complete [5]. Notably, rubric design is complex and highly time-intensive; for instance, creating the rubric for T1LegalMDS required over 10 hours of expert effort. The selection of these tasks adhere to specific criteria: (1) it is possible to create well-defined rubric for consistent and objective evaluation; (2) it requires domain expertise for both task resolution and assessment; and (3) it is grounded in real-world expert workflows. Addition task selection details are discussed in Appendix A and Appendix B. Generally, each sample in ExpertLongBench includes the following key elements:
  • Task Input: The context of the task. For example, the inputs for T2LegalSFG consist of relevant transcripts used to generate a Statement of Fact, with each sample on average having 14 documents totaling 70k input tokens.
  • Human Reference: Each task includes human-authored references in natural language. Details can be referred to Appendix B.
  • Checklist-mapped Reference: A checklist-based, fine-grained evaluation rubric is designed for each task through collaboration with domain experts and by referencing established resources. A corresponding checklist-mapped reference is constructed for every sample, as detailed in §3.2.
Tab. 1 summarizes key statistics of our benchmark. In total, our benchmark contains 1,050 samples, with an average input length and human reference length of 36,204 and 851 tokens, respectively. Specifically, T2LegalSFG has the highest average input and output length. For each task, we describe the task definition, significance, data acquisition and preprocessing, representative examples, evaluation rubric, and the construction of checklist-mapped references in Appendix B.
3.2 Benchmark Construction Procedure
Task Sources.
ExpertLongBench draws data from two primary sources: 6 newly curated task data and 5 adapted existing task data as shown in Tab. 1. The first source includes tasks newly developed by the authors specifically for this benchmark, designed to address gaps in existing resources and better capture real-world expert challenges. Tasks are selected by discussing with domain researchers and practitioners, who also contributed to rubric design. In parallel, we also look into existing datasets and corpora that fit the goal of accessing the capabilities of LLMs on expert-level tasks. To adapt these for our benchmark, we carefully select representative and challenging samples from the original datasets with details and examples shown in Appendix B. Moreover, we design a checklist-based rubric for the tasks and create checklist-mapped references for the samples.
Expert-guided Rubric Design.
For each task, we design a checklist-based fine-grained evaluation rubric to assess model performance. Our rubric design follows two complementary approaches: 1) Expert-guided design: we collaborate closely with domain experts to co-design evaluation criteria that reflect professional standards and practical requirements and needs. 2) Protocol-refinement design: for tasks with established evaluation protocols guided by experts, we design fine-grained criteria by refining and expanding upon existing standards, ensuring more granular and systematic assessment of model performance. An example rubric showcasing a subset of its checklist items is illustrated in Fig. 1. In appendix B, we show all checklist items for each task. Notably, these rubrics are written by experts, ensuring high-quality criteria that existing LLMs cannot yet replicate (refer to Appendix LABEL:ap:add_discussion). In future work, we will explore methods for automatically generating high-quality checklists. Additional examples along with detailed rubric designs for each task are in the Evaluation Rubric section for each task in Appendix B.
Sample Selection
We establish criteria for further refining the sample pool in tasks with an initial sample size exceeding 100, ultimately selecting 100 representative samples for each task. Our selection criteria focus on two key aspects: diversity and difficulty. For diversity, the select samples encompass a broad range of variations. For example, we select samples covering a wide range of sectors and companies for T10FinanceESG and prioritize the diversity of diagnosis types for T9MedicalDR. For difficulty, we identify key factors influencing difficulty for each task. For instance, we collaborate with legal researchers to determine factors such as document length, number of appeals, number of complaints, and number of dockets for T1LegalMDS. We applied these criteria across most tasks. Details are provided in the Appendix B.
Checklist-mapped Reference Creation
To better evaluate models’ capability for long-form generation on expert-level tasks according to the designed fine-grained rubric, we construct checklist-mapped references by extracting the content corresponding to each checklist item from the human reference using GPT-4o. Exceptions include T3MaterialSEG and T10FinanceESG, where the checklist-based references are constructed during the data collection process (T3) or the human references are well-structured and can simply process to structured checklist-mapped reference (T10). Example checklist-mapped references for T1LegalMDS are shown in Fig. 1 and the checklist-mapped reference subsections in Appendix B.
Specifically, we adopt a role-playing strategy by prompting GPT-4o to extract all relevant information for each checklist item as comprehensively as possible from the reference. If no such information is present, the model is instructed to return “N/A”. Example extraction prompts are illustrated in the checklist-mapped reference subsections in Appendix B. Throughout this paper, we will use the following notation: the checklist for a given task is denoted as $\{c_{i}\}_{i=1}^{n}$, where $n$ is the number of checklist items associated with the specific task. To operationalize this extraction, we design instructions for each checklist item and use GPT-4o to extract the checklist information.
Among the tasks where mapping references to checklists is straightforward, we selected two tasks for both human evaluation and LLM-as-judge evaluation to ensure quality. Evaluation results show that the mapped references achieve over 90% faithfulness and coverage. Detailed information is provided in §4.2. For other tasks, human inspection was employed to ensure the mapping quality. Additional details can be found in Appendix B.
4 Checklist-based Performance Assessment using Clear
In this section, we introduce Clear for expert-aligned evaluation of model performance on expert-level tasks. As shown in Fig. 1, given a model output, our framework maps its information to items in the checklist derived from the expert-design rubric, and assesses the quality of the model output by comparing the checklist-mapped model output against the checklist-mapped human reference (§4.1). We quantitatively justify key design choices in our evaluation pipeline, including the selection of the checklist-mapper and the judge for checklist comparison (§4.2).
4.1 Evaluation Process
Generating Checklist-Mapped Model Responses.
We follow the same procedure described in §3.2 to extract checklists from model responses. However, instead of using GPT-4o, which can be significantly more expensive while extracting checklists from all models, we opt for the open-source model Qwen2.5-72B as the checklist mapper considering its availability of model weights and decent extraction performance, which will be validated in §4.2.
Assessing Response Quality using Checklists.
To evaluate checklist-mapped responses, we assess the degree to which the checklist-mapped model response aligns with that of the reference. To this end, we use GPT-4o within an LLM-as-a-judge paradigm [24, 25, 26], adapting the reference-based scoring methodology [27] to evaluate the semantic alignment between each checklist item in a model response and the corresponding information in the reference for every sample. For instance, consider a checklist item $c$ with corresponding information in the model response and reference denoted as $H(c)$ and $R(c)$, respectively. The LLM judge assigns a binary score to each checklist item by evaluating whether the semantic content of $R(c)$ is contained within $H(c)$. The prompt for the judge is included in Appendix  LABEL:ap:eval-procedure. Despite suggestions to employ multiple models from different families as judges to mitigate biases [27], we use only GPT-4o, as it saves cost and shows high agreement with candidate judge models, as discussed in §4.2.
After performing item-level assessment for a checklist-mapped model response, we define its checklist precision (checklist recall) as the fraction of checklist items whose model response (reference) is semantically contained within the reference (model response) and accuracy as the fraction of checklist items whose model response and reference mutually contain each other. The sample-level checklist F1 score is the harmonic mean of the checklist precision and checklist recall, and we obtain the task-level performance by averaging the sample-level metrics. [^2] For simplicity, in the rest of the paper, we use precision, recall, and F1 score to indicate these checklist-specific metrics.
4.2 Evaluation Component Validation
Model Selection for Checklist Mapper. To enable accessible and low-cost checklist mapping, we evaluate open-weight models—Llama-3.3-70B-Instruct, Mistral-Large-Instruct, and Qwen2.5-72B —using reference checklists extracted by GPT-4o. We, first, validate the quality of these reference checklists through human and automated evaluations on tasks T1 and T6, confirming over 90% faithfulness and coverage. We selected these two tasks due to their challenging long contexts and extended output requirements involving over $25$ checklist items. To identify a suitable open-weight mapper, we evaluated tasks T1, T6, T7, and T8 which spans diverse domains and configurations of input / output length. Qwen2.5-72B achieved the highest average performance with an average F1-score of 90.1%, demonstrating its applicability for accurate checklist mapping. See Appendix LABEL:app:judge_analysis for details on above analyses.
Model Selection for Checklist Evaluation.
To support the choice of solely using GPT-4o for evaluating the checklist, we measured the alignment between the annotations produced by GPT-4o and Gemini-2.0-Flash on the aforementioned data used for examining checklist mapping by open-weight models. The observed Cohen’s Kappa scores were $0.81$ for T1, $0.87$ for T6, $0.89$ for T7, and $0.85$ for T8, indicating a near-perfect level of inter-annotator agreement. These high agreement scores validate the choice of using GPT-4o annotations exclusively for the final evaluations.
5 Experiments
We evaluate 11 models from 3 open-weight families (Llama [28] 8B, 70B; Mistral 12B, 123B; and Qwen [29] 7B, 72B) and 3 proprietary families (GPT-4o, GPT-4o-mini; Claude-3.5-Haiku, Claude-3.7-Sonnet; and Gemini-2.0-Flash  [^3] We include only Gemini-2.0-Flash, as the Gemini-2.0 series lacks a variant comparable in size to GPT-4o or Claude-3.7-Sonnet. Gemini-2.5 models are excluded due to frequent empty outputs without explanation.) on ExpertLongBench. For open-weight models, we use their instruction fine-tuned and RLHF variants. Model details are provided in Appendix LABEL:app:experiment_details.
Table 2: Evaluating LLMs on ExpertLongBench (scaled to 0–100) using F1 scores. Models are sorted by average performance and the best performing model on each task is bolded. Model ranking is indicated by the color of the cell, with green (best) to white (worst).
Model T1 T2 T3 T4 T5 T6 T7 T8 T9 T10 T11 Avg
Gemini-2.0-Flash 13.1 7.9 19.5 49.3 30.2 20.8 35.6 9.0 46.0 36.7 26.5 26.8
GPT-4o 13.2 6.2 15.2 56.9 29.9 25.3 34.5 10.1 35.2 35.7 29.2 26.5
GPT-4o-mini 16.5 5.9 15.2 49.2 25.0 25.5 33.9 10.3 29.5 42.2 34.2 26.1
Llama-3.3-70B-Instruct 12.1 4.9 16.8 49.5 15.9 20.2 33.4 8.3 32.8 42.6 33.8 24.6
Mistral-Large-Instruct 9.3 4.0 17.9 51.8 19.2 24.1 33.6 9.0 19.5 39.4 36.2 24.0
Qwen2.5-72B 12.3 4.1 17.4 51.0 10.8 21.2 32.6 9.3 35.7 33.3 33.9 23.8
Mistral-Nemo-Instruct 4.5 1.5 16.1 40.2 27.6 24.9 33.8 8.5 23.3 50.8 30.1 23.8
Llama3.1-8B-Instruct 9.5 3.3 20.7 48.2 18.4 23.9 32.4 6.2 25.1 40.4 29.1 23.4
Claude-3.7-Sonnet 11.5 0.9 18.1 30.4 35.0 26.1 36.1 9.2 21.4 33.1 33.5 23.2
Qwen2.5-7B 10.7 4.0 16.5 42.5 12.7 22.0 34.8 8.7 15.1 35.0 23.8 20.5
Claude-3.5-Haiku 2.8 1.1 22.1 30.3 9.7 10.9 33.4 9.2 18.1 40.8 34.1 19.3
Tab. 2 presents the performance of the evaluated models, measured with the F1 score. The models are sorted by their average F1 score across all tasks. Accuracy, precision, and recall are reported in LABEL:tab:main_accuracy_results, LABEL:tab:main_precision_results, and LABEL:tab:main_recall_results. The tasks in ExpertLongBench pose significant challenges to existing LLMs. Notably, Gemini-2.0-Flash, the top-performing model in our evaluation, achieves an average F1 score of only 26.8, underscoring the difficulty of the benchmark. Among all tasks, T2 proves to be the most challenging, with all models scoring an F1 below 8. In addition to the complex nature of legal argumentation, T2 involves longer inputs and necessitates longer outputs, thereby further testing the long-context capabilities of the models.
Scaling does not consistently improve performance across tasks.
We observe that within the same model family, the larger model outperforms their smaller counterpart in terms of average performance. However, this superiority is not consistent across all individual tasks. For example, on task T4, Mistral-Large-Instruct ranks 2nd while Mistral-Nemo-Instruct ranks 9th. Conversely, on task T10, Mistral-Large-Instruct ranks 6th, whereas Mistral-Nemo-Instruct achieves the best performance across all models. In fact, no single model consistently outperforms its smaller variant from the same family across all tasks. This inconsistency may suggest that the pre-training curricula of existing LLMs might not uniformly emphasize or adequately represent the domain-specific tasks, especially when scaling for larger models, potentially leading to an imbalance in their capabilities across different tasks.
Proprietary models are not always superior.
Figure 2: F1 score vs. coverage of checklist items. The coverage of checklist items is the percentage of checklist items that are covered in the generated text regardless their correctness.
Proprietary models, which are expected to be larger and more capable of handling diverse tasks, are not always better than open-weight models. Specifically, Claude is less designed for expert-level workflows. Compared with other families, Claude-3.7-Sonnet only outperforms Qwen2.5-7B, and Claude-3.5-Haiku yields the lowest overall score. Among open-weight models, Qwen generally demonstrates the weakest performance. Both Qwen models have a 32K token context length, shorter than other models which offer 128K or longer. This discrepancy in context length leads to a significant performance disparity for Qwen on T10, which requires processing an average of 64K input tokens, where the performance of Qwen2.5-72B trails the leading model by 17.5.
High coverage of checklist items does not imply high output quality.
We further investigate the percentage of checklist items covered in the generated text, irrespective of correctness, and its correlation with output quality, as measured by the F1 score. Fig. 2 illustrates the relationship between checklist item coverage and the F1 score, aggregated across all experimented models and tasks, as well as presented separately for task T10. A positive correlation between coverage and F1 score is observed for most tasks, resulting in an overall positive correlation when averaged. However, the correlation coefficient remains low—models frequently obtaining high coverage scores while concurrently exhibiting low F1 scores, suggesting their ability to produce content that seemingly adheres to domain-specific standards but incorporates incorrect information. Such behavior presents a risk, as the output might mislead users into perceiving the generated content as reliable due to its apparent completeness, despite its underlying inaccuracies. The issue is notably more pronounced for T10 (ESG report summarization), where coverage exhibits a negative correlation with output quality. The structure of ESG reports provides hints for elements that should be covered in the output. However, due to the volume of information, models cannot precisely identify which information is most critical to stakeholders. Further main results and detailed outcomes for each task can be found in Appendix LABEL:ap:additional_results.
6 Towards Reproducible and Low-Cost Evaluation
While GPT-4o demonstrates strong performance in evaluating checklist-aligned model responses [30, 31, 32, 33, 34], there are three primary concerns associated with the use of closed models: (a) they are typically accessible only through paid APIs, with costs that can escalate rapidly in large-scale experiments involving multiple queries or extended outputs; (b) their deployment on external servers raises data privacy and security concerns, particularly in sensitive domains such as healthcare, law, and education; and (c) proprietary models are subject to unannounced updates or version changes, which can compromise the reproducibility of results across studies or even within the same research group over time.
Therefore, the primary objective of this section is to investigate whether the judge component in our evaluation pipeline can be substituted with open-weight alternatives to promote a more accessible and reproducible research environment. To this end, we evaluate various open-weight models and their combinations as judge models by measuring how well their accuracy scores correlate with those from GPT-4o, which serves as the ground truth due to its strong domain-level performance. For model combinations involving multiple judge models, we explore two pooling strategies [35, 27] to compute an aggregate score for each sample: (a) Mean Pooling, where the final score is the average of the scores assigned by each constituent judge; and (b) Majority Pooling, where the final score corresponds to the mode of the individual scores. In cases where multiple modes exist, we resolve the tie by averaging their values.
Figure 3: Correlation of different model combinations with GPT-4o judgments averaged over all the tasks.
The results presented in LABEL:tab:main_correlation_results show the correlation among the computed scores of various model variants. In this table, the category labeled Small combines scores from Llama3.1-8B-Instruct, Mistral-Nemo-Instruct, and Mistral-Large-Instruct. Similarly, the category Large includes Llama-3.3-70B-Instruct, Mistral-Large-Instruct, and Qwen2.5-72B. The category All-small encompasses models from Small and GPT-4o-mini.
Best performing single model judge.
From Fig. 3, we find that most judge models exhibit a high degree of correlation with GPT-4o with Qwen2.5-72B yielding the highest average correlation among all single judges. This indicates that a single model can reliably assess checklists while significantly maintaining computational overhead to a low value.
Combining models results in better correlation than its individual counterparts.
Additionally, combining open-source models enhances alignment with GPT-4o. For instance, while Llama3.1-8B-Instruct, Mistral-Nemo-Instruct, and Qwen2.5-7B individually achieve average correlations of 0.65, 0.80, and 0.78, respectively, majority pooling their outputs raises the correlation to 0.82. Moreover, in most model combinations—excluding large-open-source-models—majority pooling consistently outperforms mean pooling in terms of correlation. Since majority pooling is more effective for estimating binary variables [27], and our checklist scoring relies on binary labels, it leads to more accurate overall score assignment.
Tasks T1 and T2 exhibit the lowest average correlations as shown in LABEL:tab:main_correlation_results and weakest model performance referring to Tab. 2, indicating that higher task complexity may hinder reliable evaluation.
Figure 4: Regression analysis between the average performance along each task and the average correlation of judgments with GPT-4o assignments
Connection between task complexity and judgment correlation with GPT-4o.
A regression analysis of the relationship between task complexity and average judgment correlation with GPT-4o (see Fig. 4) reveals a strong correlation and a moderately high $R^{2}$ value, indicating that the regression model explains a substantial portion of the variance. This can be used to decide the scale of model for evaluating a particular task. For instance, smaller models like Qwen2.5-7B and Mistral-Nemo-Instruct align well with GPT-4o for less complex tasks. For more challenging tasks such as T1 and T2, larger models like Qwen2.5-72B are recommended.
7 Skill Decomposition Analysis
We provide a detailed skill-level and difficulty-level analysis of model performance on ExpertLongBench for a better understanding of where current models excel and where they fall short. Our analysis is grounded in a fine-grained examination of each task’s checklist items, identifying the specific skills required to fulfill each item at corresponding difficulty levels. Detailed explanations of the skills, difficulty levels, and item-level skill and difficulty mapping are provided in Appendix LABEL:app:skill_analysis.
We examine model performance across varying knowledge levels including Below College, College, and Graduate, and reasoning difficulty including Low (Knowledge Memorization), Medium (Knowledge Understanding), High (Knowledge Applying), Very High (Knowledge Creating) [36, 37]. We present our key findings as follows.
(a) Model performance across various levels of knowledge difficulty.
(b) Model performance across various levels of reasoning complexity.
Figure 5: Model performance across various difficulty levels.
Knowledge complexity and reasoning difficulty are major barriers.
The consistent performance drop at the Graduate-level knowledge as shown in Fig. 5(a) highlights a substantial challenge for current models in mastering expert-level knowledge. The clear performance decline with higher reasoning difficulty as shown in Fig. 5(b) emphasizes the importance of assessing models beyond knowledge memorization and understanding, focusing on their ability to perform complex reasoning. Most models perform worse on low reasoning difficulty level (knowledge memorization) than medium reasoning difficulty level (knowledge understanding) because the expert-level knowledge in ExpertLongBench may be long-tail knowledge that rarely appears during training, limiting memorization performance, while the model can still leverage its reasoning and generalization abilities to approximate understanding [38]. Notably, ExpertLongBench is specifically designed to assess more advanced knowledge processing skills.
Importance of diverse difficulty levels in benchmarking.
Models show divergent strengths. For instance, Claude-3.7-Sonnet ranks first in the low reasoning level (knowledge memorization) but ranks relatively low in the very high reasoning level (knowledge creation) (Fig. 5(b)). These results emphasize the value of benchmark with diverse difficulty levels, which captures performance differences across varying knowledge and reasoning levels, revealing model strengths and weaknesses that might be overlooked in general benchmarks. Notably, ExpertLongBench is meticulously designed to provide a comprehensive evaluation across diverse difficulty levels, and we recommend the community to leverage ExpertLongBench for future model evaluation. Analysis results on more general skills and additional details are in Appendix LABEL:app:skill_analysis.
8 Conclusion
We introduce ExpertLongBench, a multi-domain expert-level benchmark. ExpertLongBench comprises 1050 samples, spanning 11 tasks across 9 distinct domains. These tasks, originating from realistic expert applications rather than typical question-answering scenarios, demand long-form outputs that adhere to domain-specific standards. Each task is accompanied by an expert-designed and validated rubric, which serve as evaluation guidelines. To evaluate long-form model outputs on these expert tasks, we design Clear, an evaluation framework that maps model outputs and references into checklists derived from the rubric and subsequently performs an item-by-item comparison of these checklists. This process yields a fine-grained, grounded evaluation that aligns with domain-specific requirements. We conduct experiments with 11 LLMs, revealing that ExpertLongBench pose a significant challenge for current LLMs. While these models can generate content that superficially matches required aspects, they frequently lack accuracy. Further analysis of Clear demonstrates the feasibility of substituting proprietary models with open-weight alternatives for the roles of checklist mapper and evaluator, facilitating low-cost, scalable benchmarking.
Acknowledgments
This project is partially supported by the NVIDIA Academic Grant Program. This work is supported in part by the National Science Foundation through grants 2046016 and 2302564. Shuyang Cao is supported by a Bloomberg Data Science Ph.D. Fellowship. We would like to thank the following domain experts for engaging in thoughtful discussions and sharing their perspectives during the development of this work. Their input helped inform our understanding across a range of specialized fields. For discussions related to ESG, we thank Daniel Cremin, Lauren Yeung, Rumi Mahmood, and Umar Ashfaq from MSCI Inc., as well as Mohammed Ombadi. For health-related topics, we thank Sharon Kardia. We are also grateful to Sabina Tomkins and Derek Van Berkel for discussions on environment and sustainability, Charles Brooks for conversations on chemistry, and Wenhao Sun for insights related to material science. We appreciate the time and perspectives shared by Patrick Barry in the legal domain. For education, we thank Xu Wang, Anne Gere, Xinyi Lu, and Jason Godfrey for their helpful discussions and references. We would also like to thank Aditya Tambe, Evan Wang, and Kaelyn Lin for their assistance on the project. Finally, we appreciate the discussions and suggestions from Michigan LAUNCH Lab members.
Appendix A Benchmark Description
Task Selection Considerations.
We follow a set of rigorous task selection criteria designed to meet our objective to accessing LLMs’ capability to solve expert-level tasks:
  1. 1.
    Human reference outputs: We only include tasks for which reliable human reference outputs are either directly available or can be accurately derived.
  2. 2.
    Long-form outputs: We focus on tasks with open-ended and long-form outputs, avoiding simple information extraction or classification tasks that lack substantial natural language output.
  3. 3.
    Objectivity: All selected tasks must be objective, allowing multiple annotators to reasonably agree on the correctness of a response.
  4. 4.
    Domain expertise requirement: Tasks must also require domain expertise, both for task completion and for evaluating the outputs.
  5. 5.
    Real-world importance and impact: We emphasize tasks with clear importance and real-world impact, particularly those that reduce expert workload or improve efficiency, accessibility, or scalability in professional workflows.
  6. 6.
    Targeting real-world users: We ensure all tasks are designed to target realistic problem-solving scenarios.
For tasks adapted from existing datasets, we apply two additional filters:
  1. 1.
    Accessibility: by selecting only publicly available datasets
  2. 2.
    Recency: giving preference to datasets released recently to minimize potential contamination from model training data.
Key Elements of Samples.
Each sample in ExpertLongBench includes the following key elements:
  • Task input: The context of the task. For example, the inputs for T2LegalSFG consist of relevant transcripts used to generate a Statement of Fact, with each sample on average having 14 documents totaling 70k input tokens.
  • Human reference: Each task includes human-authored references in natural language. Exceptions include T3MaterialSEG and T9MedicalDR, for which checklist-based references are constructed during the data collection process.
  • Checklist-mapped reference: A checklist-based, fine-grained evaluation rubric is designed for each task through collaboration with domain experts and by referencing established resources. A corresponding checklist-mapped reference is constructed for every sample, as detailed in §3.2 in the main paper.
Checklist-mapped Reference Creation.
To ensure the quality of our LLM-based checklist-mapped references, we adopt the following two strategies: For tasks where mapping references to checklists is straightforward (e.g., T1, T2, T5, T6, T7, T8, T10), we randomly select two tasks (T1, T6) and conducted human evaluation and LLM-as-judges evaluation for ensuring the quality. Details are shown in §4.2 in the main paper. For other tasks (e.g., T4, T11), we perform manual quality assurance on the model’s extracted checklist-mapped references for all samples. This process involves human reviewers inspecting and correcting any extraction errors. The remaining tasks (e.g., T3, T9) directly produce the checklist-mapped reference and no prompting is needed.
Appendix B Task Description
B.1 T1LegalMDS
B.1.1 Task Definition
Legal case summarization involves generating concise, coherent, and informative summaries from multiple legal documents pertaining to a single case. These documents can include a wide range of materials such as complaints, appeals, court opinions, motions, filings, judgments, and other related legal records. The objective is to summarize critical aspects of a case into a summary that is accessible to a broad audience, including legal professionals such as policymakers, advocates, researchers, educators, and students, as well as the general public. When using LLMs to conduct the legal case summarization task, the input is typically a collection of relevant legal documents associated with a specific case and the output is a case summary that encapsulates the essential information from the original documents.
B.1.2 Task Significance
Legal case summarization is a foundational task for improving access to justice, supporting legal research, and facilitating informed decision-making across the legal system. Many real-world cases involve large volumes of interrelated documents. These include complaints, motions, rulings, and settlement agreements, which must be synthesized into concise, coherent summaries. Automatic summarization addresses this issue by condensing lengthy texts into concise summaries, enabling users to quickly grasp essential details without reading the entire document [5]. This automation enhances accessibility by providing clear summaries for lawyers, scholars, and the general public, improves efficiency by assisting legal professionals in quickly understanding case details, and supports research by highlighting crucial aspects of cases. By reducing the time and effort required to comprehend extensive legal materials from legal experts, automatically generated summaries streamlines legal workflows and aids in informed decision-making. This challenge becomes even more pronounced in domains where litigation is prolonged, procedurally complex, and socially impactful.
One such domain is civil rights litigation, which has played a pivotal role in reforming public institutions and shaping policy in areas such as education, policing, incarceration, and disability rights. These cases often span years and generate hundreds of filings, yet most remain inaccessible to the broader public, scholars, and even practitioners. The Civil Rights Litigation Clearinghouse [^4] https://clearinghouse.net/, a leading repository in this space, curates extensive documentation from landmark and ongoing civil rights cases—many of which have never resulted in published judicial opinions. As a result, key legal developments are often buried in unstructured documents that are difficult to access and interpret without expert knowledge. Multi-document case summarization helps bridge the accessibility gap. It enables non-experts to understand legal developments, allows legal professionals to quickly grasp the structure and outcome of complex cases, supports advocacy and policymaking by surfacing systemic patterns, and reduces the burden of reviewing lengthy legal documents, giving researchers easier access to representative case content. We collect data from Clearinghouse, with further details provided in Section B.1.3.
A related work, Multi-LexSum [5], also investigates the task of multi-document legal case summarization in civil rights litigation domain using data from Clearinghouse. However, their study relies on automatic evaluation metrics such as ROUGE-{1,2,L} [16] and BERTScore [18]. In contrast, our work constructs a new dataset by collecting cases data from the Clearinghouse and introducing clear selection criteria with legal researchers grounded in both diversity and difficulty, where diversity refers to the range of case topics and difficulty is based on structural factors such as such as the number of documents, length, and complexity of legal proceedings. Furthermore, we develop a fine-grained, checklist-based evaluation rubric in collaboration with a legal researcher to ensure comprehensive coverage of key case elements. To assess model performance more reliably, we design a structured evaluation pipeline leveraging LLMs as evaluators, enabling nuanced analysis beyond surface-level lexical similarity.
B.1.3 Data Acquisition and Preprocessing
We collected multi-document legal case summarization data with 1,393 legal case samples from a widely recognized legal repository, Clearinghouse. We obtained the case data through the official Clearinghouse API [^5] https://clearinghouse.net/api, which provides structured access to metadata and associated documents for civil rights litigation cases. For documents available in PDF format, we used the ocrmypdf [^6] https://github.com/ocrmypdf/OCRmyPDF tool to extract text from scanned legal filings, ensuring complete retrieval of document content. Each sample includes multiple case documents. Alongside these inputs, the dataset provides expert-crafted, gold-standard summaries that distill the key facts, legal principles, and outcomes of each case. These summaries are produced by trained legal experts—including lawyers and law students—who follow detailed annotation guidelines. To further ensure consistency and accuracy, each summary undergoes an additional round of expert review. We then selected a small, high-quality and representative subset from the large-scale dataset to evaluate the ability of LLMs in generating legal summaries.
For representative data selection, we filter the data based on the following standards to select a high-quality dataset.
  • Diversity: We obtain the topic for each case on the Clearinghouse website and select cases spanning as many distinct topics as possible. The selected data covers diverse case types such as “Equal Employment”, “Prison Conditions”, “Election/Voting Rights” which contain the specific topic the legal case is dealing with.
  • Difficulty: We identify several factors influencing difficulty after discussing with legal researchers. The factors are as follows:
    • Length of the documents
    • Number of appeals: Appeals are processes that the litigant asks for the higher court review of the lower court’s decision, we obtain the number of appeals for each case from the Clearinghouse website.
    • Number of complaints: Complaints are documents that start the lawsuit. We obtain the number of complaints for each case from the Clearinghouse website.
    • Number of dockets: Dockets are logs containing a comprehensive history of the case in chronological order. We obtain the number of dockets for each case from the Clearinghouse website.
After this process, we select 454 samples with a high difficulty level.
To further ensure the quality of the gold-standard case summary, we discuss with legal researchers and identify the required checklist items that must be included in the summary. We then use GPT-4o as a judge to assess whether the human-written reference covers the information and select the cases with highest-quality human summaries, determined by the highest coverage of checklist items. The specific prompt used for accessing the reliability of human-written summary is shown in Tab. 3. Based on the judgment results, we finally select 100 samples with high-quality and long human reference summaries.
Table 3: T1LegalMDS - Prompt for checking the quality of human-written references.
Evaluate whether the given summary includes all the information listed in the checklist items. For each checklist item, provide whether the item is fully addressed (“Yes” or “No”) and a brief explanation or evidence supporting your evaluation. Finally, provide an overall result summarizing whether all checklist items are covered in the given summary. Provide the overall result after ####, eg #### Yes or #### No. Given Summary: [REQUEST] Checklist Items: Filing Date Whether it contains the Date: yes/no Cause of Action Description: e.g., a statute (e.g., 42 USC 1983) or a case (e.g., Ex Parte Young) Whether it contains the information: yes/no * yes: the summary clearly states the action information * no: do not mention the action information Statutory or Constitutional Basis for the Case Description: A case can either be based on a statute or a provision of the Constitution–i.e., a case will either claim that someone violated a statute, or violated the Constitution. For cases that have a constitutional basis, the summary should refer to the clause of the Constitution that was allegedly violated, as well as the amendment if applicable. So for example it would say “the plaintiffs alleged violations of the Fourteenth Amendment’s Equal Protection Clause,” or “the plaintiffs alleged violations of the Commerce Clause.” Whether it claim that someone violated a statute or violated the Constitution: yes/no Whether it contains the statutory bases information or constitutional bases information: yes/no * yes: contains the statutory bases or constitutional bases information * no: do not contain any statutory basis or constitutional bases information Remedy Sought Description: e.g., declaratory judgment Whether it contains the information: yes/no Who are the parties (description, not name)? Whether it contains the information: yes/no * Whether it contains the plaintiff information * Whether it contains the description of the defendants (usually based on their office/position if it’s a government official): yes/no Type of Counsel Description: type of counsel contains private, legal services, ACLU, etc. Whether it contains the information: yes/no First and Last Name of Judge Description: Form: Judge John Smith. Whether the reference includes this information and the generated result also includes it, or the reference does not include this information and the generated result also does not include it: yes/no Factual Basis of Case Description: Refers to the facts or evidence upon which the case is built. These facts are essential in the legal process and are used to support legal claims or decisions. It typically includes: 1. Details of the relevant events: For example, what happened, when it happened, where it happened, and who was involved. 2. Evidence: Physical evidence, documentary records, witness testimonies, etc., that support these facts. 3. Background information: Context or explanatory facts that provide additional understanding. In legal proceedings, the factual basis is crucial for determining the outcome of a case, as the judge or jury makes decisions based on the facts and the applicable legal principles. Whether it contains the information: yes/no
B.1.4 Illustrative Example
An illustrative example is shown in Tab. 4. All samples for T1 are publicly available [^7] https://huggingface.co/datasets/launch/ExpertLongBench.
The sample input for T1 consists of multiple legal case documents from a class-action lawsuit concerning COVID-19 conditions in immigration detention facilities. These include court filings such as status reports, motions, court orders, and objections, all focused on health and safety concerns at the Mesa Verde Detention Facility and the Yuba County Jail during the pandemic. Due to the length of the original documents, we provide only brief descriptions of each one’s purpose and topic. The human reference output is an expert-curated summary that concisely captures the key legal developments and arguments in the case. It outlines the plaintiffs’ claims regarding unsafe detention conditions, the legal basis for the lawsuit, major court decisions (e.g., temporary restraining orders and preliminary injunctions), and procedural milestones such as appeals and mediation efforts.
Additionally, Tab. 5 presents the model prompt. The prompt was carefully constructed to be high-level and general instead of providing the exact checklist items. The inputs for this prompt are the documents shown in Tab. 4.
Table 4: T1LegalMDS - A sample of case documents and the corresponding case summary.
Case documents (Sample input): Case document 1: { Case 3:20-cv-02731-VC, Document 407 — Defendants’ Joint Status Report re COVID-19 Testing at Mesa Verde; Topic: COVID-19 testing plans and a positive case at Mesa Verde Detention Facility } Case document 2: { Case 3:20-cv-02731-VC, Document 40 — Plaintiffs’ Opposition to Motion to Stay in Light of Fraihat; Topic: Argument against staying proceedings due to a related case (Fraihat v. ICE) } Case document 3: { Case 3:20-cv-02731-VC, Document 658 — Plaintiffs’ Objections to Defendants’ Medical Plan; Topic: Critique of Defendants’ COVID-19 medical care plan for Mesa Verde } Case document 4: { Case 3:20-cv-02731-VC, Document 595 — Court Order to Create Medical Plan; Topic: Court mandates improvement in COVID-19 medical response } Case document 5: { Case 3:20-cv-02731-VC, Document 922 — TRO re: COVID-19 Conditions at Yuba County Jail; Topic: Temporary Restraining Order to protect detainees at Yuba County Jail } Case summary (Human reference): COVID-19 Summary: This is a class-action lawsuit brought on April 20, 2020, by seven individuals in immigration detention at the Mesa Verde Detention Facility (MVDF) and the Yuba County Jail (YCJ), seeking immediate release from unsafe conditions of the jail in light of the global coronavirus pandemic. The court granted the request for TRO on April 29, requiring ICE to provide information and access to detainees to facilitate a process of considering bail requests. On June 9, the court granted the motion for preliminary injunction and ordered the defendants to maintain the status quo while the case was pending. On August 5, the plaintiffs sought a TRO, claiming that the defendants’ actions were insufficient and it was granted the next day. After the parties agreed to implement testing and other public safety protocols, the defendants filed a motion to dismiss and the plaintiffs responded with an amended class-action complaint. The court granted a second preliminary injunction on December 3, 2020, and a preliminary injunction related to a YCJ outbreak on January 6, 2021. The defendants appealed to the Ninth Circuit and the parties entered into mediation.
Table 5: T1LegalMDS - Model prompt.
Generate a clear and legally precise summary of a multiple-document legal case. Focus on capturing key facts, procedural history, and significant rulings in a way that is easy to understand. Provide enough detail to convey the case’s development and outcome without being excessively long or overly detailed. These are the case documents:
B.1.5 Evaluation Rubric
We closely collaborated with legal researchers to design a checklist-based evaluation rubric to assess the helpfulness of legal case summaries. The process of designing, refining, and finalizing the rubric takes roughly 11 hours. This rubric comprises of 26 checklist items that capture key aspects to consider during the evaluation process. Some items are marked as “if applicable”, meaning they should only be considered when the case pertains to the corresponding checklist criteria. This terminology will be used in other task checklists as well. The detailed checklist items are listed as follows:
  1. 1.
    Filing Date
  2. 2.
    Class Action or Individual Plaintiffs? (if applicable): If there are class action plaintiffs the summary should say it’s a class action; if there are individual plaintiffs it can just describe the plaintiffs. For example, use specific terms like “The city” or “The parents” rather than general terms like “The defendant” or “The plaintiffs.”
  3. 3.
    Cause of Action: e.g., a statute (e.g., 42 USC 1983) or a case (e.g., Ex Parte Young)
  4. 4.
    Statutory or Constitutional Basis for the Case: A case can either be based on a statute or a provision of the Constitution—i.e., a case will either claim that someone violated a statute, or violated the Constitution. For cases that have a constitutional basis, the summary should refer to the clause of the Constitution that was allegedly violated, as well as the amendment if applicable. For example it would say “the plaintiffs alleged violations of the Fourteenth Amendment’s Equal Protection Clause,” or “the plaintiffs alleged violations of the Commerce Clause.”
  5. 5.
    Remedy Sought: e.g., declaratory judgment
  6. 6.
    Who are the parties (description, not name)?
  7. 7.
    Type of Counsel: type of counsel contains private, legal services, ACLU, etc.
  8. 8.
    Consolidated Cases Noted (if applicable)
  9. 9.
    Related Cases listed by their case code number (if applicable)
  10. 10.
    Note important filings (if applicable): Note important filings including motions for temporary restraining orders or preliminary injunctions, motions to dismiss, motions for summary judgment, etc.
  11. 11.
    All reported opinions cited with shortened Bluebook citation (if applicable): For example, the summary could write “2020 WL 4218003” after the paragraph in which it discusses that opinion. The summary does not need to include the case name, court, or date unless helpful—such as when the summary cites an opinion from a different case.
  12. 12.
    First and Last Name of Judge: Form: Judge John Smith. Find judge’s first names at http://www.fjc.gov/public/home.nsf/hisj
  13. 13.
    Significant Terms of Decrees (if applicable): Significant terms means the substance of the decree or settlement. In a decree, the judge orders the defendants to do something; in a settlement, the defendants agree to do something. The significant terms would be what the defendants are ordered/agree to do.
  14. 14.
    Dates of All Decrees (if applicable)
  15. 15.
    How long decrees will last (if applicable)
  16. 16.
    Significant Terms of Settlement (if applicable)
  17. 17.
    Date of settlement (if applicable)
  18. 18.
    How long settlement will last (if applicable)
  19. 19.
    Whether the settlement is court-enforced or not (if applicable)
  20. 20.
    Was there a monitor? Note the name of the monitor (if applicable)
  21. 21.
    Monitor’s Reports (if applicable): A monitor’s report explains whether a defendant is complying with a court order, so people want to know which terms of the order are being complied with. For example, from this case: In February 2016, the monitor filed her first semi-annual report. The report stated that the payment to the plaintiffs and plaintiffs counsel was made; the requirement of hiring ADA Coordinators was nearly compliant; videophone installation is apparently compliant; free access to videophone, provision of qualified interpreters for unscheduled medical emergencies, and provision of qualified interpreters for disciplinary hearings were unclear; informational materials were partially compliant; the routine and situational reporting were difficult or partially noncompliant; and training was noncompliant.
  22. 22.
    Appeal (if applicable)
  23. 23.
    Trials (if applicable)
  24. 24.
    Court rulings on any of the important filings (if applicable): This category corresponds with the “important filings” category—so whenever an important filing is mentioned, people also want to know what the ruling on that filing was (if there is one)—e.g., whether the judge granted or denied a motion to dismiss. Generally these filings would be: Motions to dismiss, Motions for summary judgment, Motions for a preliminary injunction or temporary restraining order, Motions for class certification, Motions for attorneys’ fees, Amended complaints–these won’t have rulings, so they should be in the “important filings” category but not the “rulings on to important filings” category, statements of interest–similar to above, there won’t be rulings on these.
  25. 25.
    Factual basis of case: Refers to the facts or evidence upon which the case is built. These facts are essential in the legal process and are used to support legal claims or decisions. It typically includes: 1. Details of the relevant events—For example, what happened, when it happened, where it happened, and who was involved. 2. Evidence – Physical evidence, documentary records, witness testimonies, etc., that support these facts. 3. Background information—Context or explanatory facts that provide additional understanding. In legal proceedings, the factual basis is crucial for determining the outcome of a case, as the judge or jury makes decisions based on the facts and the applicable legal principles.
  26. 26.
    Disputes over settlement enforcement (if applicable)
B.1.6 Checklist-mapped Reference
We created checklist-mapped references for the human reference and the model output based on our 26-item checklist.
For the human reference, we used GPT-4o to extract answers to each checklist item individually. Each prompt targeted a single item and followed a structured format. To improve contextual grounding, we used role-playing to frame the model as an assistant to a lawyer. If no related information was mentioned in the legal case summary, the model should return “N/A” to reduce the possibility of hallucination. An example is shown in Tab. 6.
For the model output, we use Qwen2.5-72B to create the checklist-mapped reference. Our decision to use a different model is supported in §4.1. To reduce cost while maintaining quality, we grouped the checklist items rather than prompting for them individually. We divide the 26 items into five groups, each with four to six checklist items, based on the average expected length of the model’s response. Each model output prompt used the same template, role-playing format as the human reference prompt. Due to space constraints, we release the full prompt set in our public GitHub repository.
The checklist-mapped reference of the example in Tab. 4 is presented in Tab. 7. To assess the quality of the checklist-mapped reference, we selected an additional 30 difficult and diverse samples, different from the main dataset, and conducted human verification to examine the faithfulness of the model response. Details of this experiment can be found in LABEL:ap:validation-of-checklist.
Table 6: T1LegalMDS - Prompt for extracting checklist-mapped reference.
You are assisting a lawyer in extracting key information from a legal case summary. Given a case summary, extract the cause of action, including whether it refers to a statute (e.g., 42 USC 1983) or a case (e.g., Ex Parte Young). Extract crucial related information as completely as possible. If no related information is mentioned in the legal case summary, state “N/A” (as a string). This is the case summary:
Table 7: T1LegalMDS - Checklist-mapped reference.
1. Filing Date: 04/20/2020 2. Class Action or Individual Plaintiffs? (if applicable): Class Action 3. Cause of Action: Habeas petition under 28 U.S.C. § 2241 and declaratory action under 42 U.S.C. § 2201-02 4. Statutory or Constitutional Basis for the Case: Violations of the Administrative Procedure Act, the First Amendment, and the Fifth Amendment’s Due Process Clause 5. Remedy Sought: Declaratory judgment, injunctive relief, release of detainees, prohibition of solitary confinement for social distancing, attorney fees 6. Who are the parties (description, not name)?: Plaintiff: Seven noncitizens detained at Mesa Verde Detention Facility and Yuba County Jail, represented by the ACLU of Northern and Southern California, the San Francisco Public Defender’s Office, Lawyers’ Committee for Civil Rights, and private attorneys; Defendants: U.S. Immigration and Customs Enforcement (ICE) and GEO Group, the ICE contractor operating the facility 7. Type of Counsel: ACLU of Northern California and Southern California, San Francisco Public Defender’s Office, Lawyers’ Committee for Civil Rights, and private attorneys 8. Consolidated Cases Noted (if applicable): Yes, appeals were consolidated under Case Nos. 21-15195 and 21-15197 9. Related Cases listed by their case code number (if applicable): 21-15195; 21-15197 10. Note important filings (if applicable): Temporary Restraining Orders (April 29, August 6, December 23, 2020); Preliminary Injunctions (June 9, December 3, 2020; January 6, 2021); Motion for Class Certification (April 2020); Motion to Stay (April 25, 2020); Motion to Dismiss (September 21, 2020 and January 19, 2021); Amended Complaint (October 30, 2020); Motion to Modify Preliminary Injunction (March 27 and June 1, 2021); Motion for Final Settlement Approval (May 27, 2022). 11. All reported opinions cited with shortened Bluebook citation (if applicable): 2020 WL 2059848; 2020 WL 3055449; 2020 WL 4554646 12. First and Last Name of Judge: Vince Chhabria 13.
B.2 T2LegalSFG
B.2.1 Task Definition
A Statement of Fact (SOF) is a document used in legal cases to persuade the judge of a certain viewpoint. It outlines the events and circumstances leading up to a legal dispute in an objective manner, while also incorporating persuasive language when appropriate [39]. The objective is to generate a comprehensive account of the case, based on courtroom transcripts, that accurately reflects the its development. The input includes transcripts from multiple proceedings, such as preliminary trials, adjudication hearings, and termination trials. The output should present all relevant facts, direct quotations, and procedural details with proper attribution and clear alignment to legal standards. It must stand alone as a complete summary, without requiring reference to the original transcripts.
B.2.2 Task Significance
SOFs play an crucial role in improving clarity, consistency, and accessibility in the documentation of court proceedings. They are a part of a legal brief, formal documents submitted to a court that outline the relevant facts, applicable laws, and legal arguments of a case. Our samples are specific to in child welfare and termination of parental rights cases. These cases often involve complex and emotional content that must be translated into reliable and grounded summaries.
Traditionally, drafting such briefs is a time-intensive and expert-level task. Attorneys must review hundreds of pages of transcripts to extract and organize key information into create an coherent narrative. This process often takes 20 to 40 hours or more, even for seasoned professionals [40]. Manual note-taking adds further difficulty - often leading to inconsistent phrasing, missed information or subjective interpretation. It also requires frequent references to transcripts, which not only slows progress, but increases the risk of delays and oversights in case processing.
Automating this process with LLMs addresses these challenges by generating detailed and structured briefs. They are especially crucial in child welfare systems, where timely and accurate information directly impacts case decisions and the well-being of children. In this task, transcripts are drawn from real juvenile court proceedings provided by our collaborators, and outputs are compared against a detailed rubric developed with experts to ensure both completeness and legal relevance.
B.2.3 Data Acquisition and Preprocessing
Domain experts provided us with 113 samples of transcripts in Word document or PDF form. We first processed all documents using a custom script to extract clean, structured text from each file, regardless of format. The script handles .docx, .doc, and .pdf files and includes fallback methods to ensure robustness across formatting inconsistencies.
For .pdf files, we first attempted direct text extraction using PyPDF2888https://pypi.org/project/PyPDF2/. If the PDF lacked embedded text (e.g., scanned images), we applied OCR using OCRmyPDF [^9] https://github.com/ocrmypdf/OCRmyPDF with deskewing, noise cleaning, and compression optimizations. This was the library recommended to us by our collaborators.
For .docx files, we used the python-docx library [^10] https://pypi.org/project/python-docx/ to extract paragraph-level content. For older .doc files, we employed a fallback approach using textract [^11] https://pypi.org/project/textract/, antiword [^12] http://www.winfield.demon.nl/, or conversion to .docx using LibreOffice in headless mode [^13] https://www.libreoffice.org/.
Each line of extracted text was then cleaned to remove extraneous whitespace and empty lines, ensuring that only content-relevant information was retained. However, page numbers were retained in the final text as the model prompt asked the LLM to generate SOFs that cited the transcript’s page numbers. Cleaned output was saved in plain .txt files, one per document. For each case, all .txt files were then concatenated into a single input string in the format: transcript 1 name: transcript 1 content, transcript 2 name: transcript 2 content … . This structured format served as the sample input.
To filter 113 transcript samples, we selected the 100 with the longest human-written reference SOFs, consistent with the reasoning in §3.2. On average, the 100 samples included 14 transcript documents, each 5K+ tokens in length, and paired with a human reference SOF also exceeding 5K tokens. This filtering ensured that our dataset contained the most information-dense and well-documented cases.
B.2.4 Illustrative Example
Due to the proprietary nature of the data, we cannot provide a specific example. [^14] For review purpose, data can be previewed via the dataset url in the submission form. However, in Tab. 8, we provide the model prompt for obtaining model outputs.
Table 8: T2LegalSFG - Model prompt.
You are an expert appellate lawyer conducting a comprehensive review of a legal case based on the attached transcripts. Your task is to create a chronological and unbiased narrative statement of facts, ensuring all key material details are accurately represented with specific transcript page citations. These are the transcripts:
B.2.5 Evaluation Rubric
We worked with three domain experts and drafted through three iterations of the checklist through discussion. It takes approximately 8 hours to design, revise, and finalize the rubric. The experts designed the checklist items and helped categorize each checklist item into general levels and provided brief definitions of some levels. It is important to note that some checklist items are conditional and depend on the responses to previous items. For instance, Items 26 and 27 correspond to different outcomes based on the response to Item 25. If the parent pled to adjudication, Item 26 should be completed and Item 27 will be “N/A”. If there was an adjudication trial instead, Item 27 should be used and Item 26 will be “N/A”. The detailed checklist items are listed below:
  • Initial incident
    1. 1.
      When DHHS first made contact
    2. 2.
      What caused DHHS to become involved
    3. 3.
      Discrepancies between DHHS’ account and the parent’s account
  • Background situation
    • We want to know any facts about the parent and children that can help the court understand their situation more favorably. Examples can include a parent having an unstable home environment growing up, a parent having an abusive partner, a child having disabilities, etc.
    1. 4.
      Challenges the parent faced when they were young
    2. 5.
      Problems the parent currently faces
    3. 6.
      Child’s special needs
  • How the child is doing
    1. 7.
      How the child(ren) is doing
    2. 8.
      Child(ren)’s development
    3. 9.
      Child(ren)’s interactions with the parent
    4. 10.
      Child(ren) talked about their feelings or experience
  • Impact of each court event
    1. 11.
      Decision of judge
    2. 12.
      New actions the parent or agency has to do
    3. 13.
      Actions the parent or agency does not have to do anymore
  • Details of service plan
    • In these cases, the court typically orders the parent to complete some requirements so they can get custody of their child back. Examples of “services” can include parenting classes, therapy, drug screens, etc.
    1. 14.
      Reasonable efforts to reunify
    2. 15.
      If the agency was not ordered to make reasonable efforts to reunify,what aggravating circumstances were present
    3. 16.
      Services the court required the parent to complete
    4. 17.
      For each service, what the service means
    5. 18.
      For each service, what objectives the parents have to meet
  • Parent progress on service plan
    1. 19.
      How the parent complied with the services
    2. 20.
      How the parent did not follow through with the services
    3. 21.
      Parent’s compliance or noncompliance with the service plan
  • Agency role in service plan
    • The child welfare agency might help the parent with compliance with things like transportation help and therapy referral. But the government can also hurt the parent with things like delays and gaps in communication.
    1. 22.
      How the agency tried to help the parent get the services they need
    2. 23.
      How the agency or government ignored a parent or undermined their efforts
  • Adjudication hearing
    1. 24.
      When the adjudication hearing was held
    2. 25.
      An adjudication trial or the parent pled to adjudication
    3. 26.
      If the parent pled to adjudication, the specific allegations the parent pled to
    4. 27.
      If there was an adjudication trial, what statutory grounds for adjudication were found
  • Permanency planning hearings
    1. 28.
      When each permanency planning hearing was held
    2. 29.
      What permanency plan was established
    3. 30.
      Court’s reasons for the chosen permanency plan
    4. 31.
      Whether DHHS was ordered to initiate termination of parental rights
    5. 32.
      If the child has been in foster care for more than 15 of the last 22 months, and the court did not order DHHS to initiate termination proceedings, the court’s reasoning
  • Termination of parental rights hearings
    1. 33.
      Statutory grounds DHHS sought termination
    2. 34.
      Statutory grounds the court found that termination was proper
    3. 35.
      Evidence offered to show statutory grounds for termination were not met
    4. 36.
      Evidence offered to show termination was in the child’s best interests
    5. 37.
      Evidence offered to show termination was not in the child’s best interests
    6. 38.
      If the child was placed with a relative, whether any parties discussed that placement weighing against termination
    7. 39.
      Evidence regarding what kind of permanency the child will have if the parent’s rights are terminated
    8. 40.
      Alternatives to termination of parental rights
  • Last sentence
    1. 41.
      Ends with the final lower-court ruling
B.2.6 Checklist-mapped Reference
The creation of the checklist-mapped reference is for T2 is similar to T1’s process in described in Section B.1.6.
However, unlike in T1, the checklist items used to construct the model output checklist-mapped reference are grouped differently. We divide the 41 items into nine groups based on: (1) the average expected length of the model’s response, and (2) the topical category of the item (e.g., Details of service plan, Impact of each court event, etc.). For instance, we grouped checklist items 1–6 together, as they all fall under the categories Initial incident and Background situation. This grouping strategy helps the model process large inputs more effectively by allowing it to focus on fewer sections at a time. The prompt structure for extracting an individual checklist item is illustrated in Tab. 9.
Due to the proprietary nature of the data, we cannot provide a specific example for the checklist-mapped reference.
Table 9: T2LegalSFG - Prompt for extracting checklist-mapped reference.
You are assisting an appellate lawyer in extracting key information from a Statement of Facts (SOF). Given a SOF, identify if it states when the DHHS first made contact with the family. Provide the extracted information. If when the DHHS first made contact with the family is not mentioned, state “N/A”.
B.3 T3MaterialSEG
B.3.1 Task Definition
Materials science research involves synthesizing new materials and understanding their properties. A synthesis recipe includes a series of steps for creating a target material, specifying the precursors and synthesis conditions used. In the task of synthesis explanation generation, the objective is to provide justifications for key decisions made in the recipe, considering factors such as the precursors’ structural motifs, reactivity, thermodynamic stability, and more.
B.3.2 Task Significance
A material can often be synthesized through multiple methods, and the choice of precursors and synthesis conditions can significantly affect the properties and yields of the resulting material [41]. Understanding the reasoning behind these choices is essential for materials scientists aiming to optimize synthesis protocols and develop new materials with desirable properties through more efficient processes. Additionally, materials discovery efforts, especially those aided by high-throughput computations [42, 43], can be accelerated by prioritizing synthesis recipes that are grounded in the most promising mechanistic insights. Per our discussion with materials science researchers, writing the explanations for a synthesis recipe from scratch would take a well-trained materials science PhD student 1-2 hours.
B.3.3 Data Acquisition and Preprocessing
Despite the availability of large-scale synthesis recipe datasets [43, 44], these curated recipes are not accompanied by explanations. Even for well-trained materials science PhDs, writing explanations for synthesis recipes from scratch is challenging and time-consuming. In contrast, given statements about the synthesis recipe, undergraduate-level chemistry students are able to distinguish between explanations and non-explanations. Therefore, we adopt a semi-automated approach to collect synthesis recipes along with corresponding explanations from existing literature.
We begin by consulting materials science researchers [^15] They are PhD candidates in materials science. to identify the key aspects that explanations should address. Since papers in the field do not always explicitly state the rationale behind synthesis choices, we begin by identifying papers that are more likely to include such explanations. To do this, we extract synthesis recipes and potential explanations using Llama-3.3-70B-Instruct with carefully designed instructions. [^16] We also experimented with in-context learning, but found that Llama-3.3-70B-Instruct often copied from the in-context examples rather than extracting content from the target paper. We then rank the papers based on the number of aspects covered. Starting from the top-ranked papers, we use GPT-4o to further extract synthesis recipes and corresponding explanations. Based on GPT-4o’s outputs, annotators with undergraduate-level chemistry knowledge verify the extracted recipes and collect accurate explanations. When explanations are absent from the paper, annotators are asked to discard the paper and move on to the next one, until the target number (50) of samples is collected.
We retrieve papers containing the keyword “solid-state synthesis” from Science [^17] https://www.science.org and Wiley [^18] https://onlinelibrary.wiley.com, as the aspects of interest are more relevant to solid-state synthesis. The text of the paper is extracted directly from the HTML version of the paper. During the automated extraction step, we generate five different outputs using GPT-4o for each paper to increase the recall of possible explanations. To improve the faithfulness of the extraction, we also prompt the model to provide the source sentences corresponding to each explanation. The prompt is shown in Tab. 10, which is based on the rubric designed in Section B.3.5. Feedback from the annotators indicates that some explanations extracted by the LLMs describe properties of the synthesized materials (e.g., crystal structure, conductivity) rather than justifying the synthesis steps themselves.
Table 10: T3MaterialSEG - Prompt for automatically extracting synthesis recipes and explanations.
[…]
B.3.4 Illustrative Example
An illustrative example is shown in Tab. 11. All samples for T3 are publicly available.
The sample input is a synthesis recipe for a complex oxide thin film. The recipe describes the starting precursors, key thermal steps, and atmospheric conditions used during synthesis. This format mirrors the language and structure typically found in materials science literature and experimental protocols. The human reference output provides explanations for why each major synthesis step and condition was chosen. It should help researchers understand the scientific rationale behind the recipe, not just its procedural steps.
Additionally, Tab. 12 presents the model prompt.
Table 11: T3MaterialSEG - A sample of a synthesis recipe and its corresponding explanations.
[…]
Table 12: T3MaterialSEG - Model prompt.
You are a materials science researcher. Given a synthesis recipe that includes the target material, selected precursors, and synthesis steps, your task is to justify the key decisions made in the recipe. This includes explaining the rationale behind the choice of precursors, reaction conditions, and processing steps, using relevant principles such as structural compatibility, chemical reactivity, and desired phase formation. Output the explanation rationales as a list of bullet points, where each bullet point contains complete sentences.
B.3.5 Evaluation Rubric
We work with a PhD student in material science, whose research focuses on solid-state synthesis. They provide us with five papers191919doi.org/10.1002/ejic.202100901, doi.org/10.1002/kin.21467, doi.org/10.1002/smll.202206248, doi.org/10.1126/sciadv.adj5431, doi.org/10.1126/sciadv.adp3309 that contain well-written explanations for synthesis recipes. We prompt GPT-4o, Claude-3.7-Sonnet, and Gemini-2.0-Flash to summarize the aspects of the explanations covered in these papers. The PhD student then reviews the aspects summarized by the three models, revises them, and adds additional key aspects that are not covered by the models. The final rubric includes six items:
  • Selection of Precursors
    1. 1.
      Structural Considerations: Justify precursor selection by explaining how the precursor’s structural motifs (e.g., coordination environments, lattice arrangement) influence the target phase formation.
    2. 2.
      Handling Precursor Reactivity: Justify precursor selection by explaining the impact of precursor reactivity on phase evolution.
    3. 3.
      Physical and Chemical Properties of Precursors: Justify precursor selection by addressing how precursor properties (e.g., particle size, morphology) influences reaction kinetics and product morphology.
  • Synthesis Conditions
    1. 4.
      Temperature and Heating Method: Justify the choice of synthesis temperature and heating method (e.g., based on thermodynamic considerations, reaction kinetics, heat transfer efficiency, or side reactions).
    2. 5.
      Atmosphere: Justify the choice of synthesis atmosphere environment (e.g., based on thermodynamic considerations, reaction kinetics, or side reactions).
    3. 6.
      Duration: Justify the choice of synthesis duration (e.g., based on reaction kinetics, phase transformation rates, or side reactions).
B.3.6 Checklist-mapped Reference
Synthesis explanation generation does not involve the checklist mapping process, as the annotation process directly produces the reference checklist. An example checklist-mapped reference is shown in the illustrative example in Tab. 11.
B.4 T4EduPAE
B.4.1 Task Definition
Assessing pedagogical alignment involves evaluating how well a large language model (LLM) can replicate effective teaching strategies, such as offering scaffolded guidance without directly revealing the answer [45, 46, 47, 1]. One possible approach is to evaluate LLMs through real-world interactions with students. However, this poses challenges, as student responses can vary unpredictably across different models [48]. Additionally, using the same group of students to test multiple LLMs introduces complications: once students are exposed to a topic in one session, their familiarity with the material can influence their performance in subsequent sessions, making it difficult to isolate the effectiveness of each system. Conversely, using different students for the same set of topics may lead to non-standardized assessments due to variations in their prior knowledge and skill levels. The objective is to assess whether LLMs can provide pedagogically sound feedback that supports student learning without directly giving away answers.
To provide a repeatable and standardized framework for evaluating the pedagogical abilities of different LLMs, we propose a task focused on assessing the effectiveness and pedagogical alignment of feedback within a tutor-student dialogue. Each task instance presents a multi-turn conversation in which a student attempts to solve a biology problem with the guidance of a tutor. The tutor employs a step-by-step problem-solving strategy, breaking the larger problem into smaller sub-problems and promoting active learning by guiding the student through each segment [46]. This process is sequential: the student addresses one sub-problem at a time before moving on to the next. The conversational context includes the full dialogue history up to the current sub-problem, concluding with the student’s latest response.
For each instance, the evaluated LLM is required to carry out the following: (1) assess the accuracy of the student’s response using three coarse-grained categories—accurate, partially accurate, and inaccurate; (2) determine whether the student is experiencing difficulty with the current sub-problem by detecting repeated errors across multiple attempts; and (3) provide constructive, pedagogically sound feedback that both diagnoses the underlying misconceptions and supports the student in resolving them and (4) initiate a transition to the next sub-problem, if applicable, once the student’s response is deemed satisfactory. When the student has made only a single error on a sub-problem, the feedback should be indirect and refrain from revealing the correct answer. Conversely, in cases of repeated mistakes, the feedback may include direct identification of the error and explicit corrective guidance, even if that entails disclosing the correct solution. This pedagogical plan is provided to each of the models being evaluated and we assess their ability in adhering to these instructions.
B.4.2 Task Significance
Given their remarkable capabilities across a wide range of language processing and knowledge-intensive reasoning tasks, LLMs hold great promise as powerful tools for delivering effective learning experiences through intelligent tutoring systems. However, their application in educational contexts faces two key limitations.
First, LLMs may produce inaccurate reasoning [49] or rely on incorrect information [50] when generating feedback, which can hinder rather than support student learning. Second, while LLMs are generally optimized to be helpful and harmless, the operational definition of “helpfulness” may not align with pedagogical goals. Specifically, a fundamental strategy in conceptual learning is to guide students through problem-solving processes without directly revealing the answer. In contrast, LLMs are often trained to prioritize immediate assistance, which may conflict with the educational objective of fostering deep understanding through indirect guidance.
To address these concerns, our task aims to evaluate several critical dimensions of an LLM’s pedagogical capacity: (1) Its effectiveness in adhering to an explicit pedagogical plan; (2) Its ability to apply domain-specific knowledge to accurately interpret student responses and determine appropriate next steps; (3) Its capacity to comprehend the conversational context and identify the current sub-problem being addressed; (4) Its skill in offering feedback that accurately identifies and corrects errors; and (5) Its ability to formulate feedback that encourages active learning through indirect guidance rather than direct answers.
B.4.3 Data Acquisition and Preprocessing
We first provide a background on the source of the data that will be used for building this task. We used the multi-turn student-tutor conversations simulated using GPT-4o by [51]. These dialogues are grounded in the socio-constructivist model of learning [52], where the tutor adopts a supportive and encouraging tone while offering feedback through Socratic questioning and indirect hints, rather than direct instruction. An illustrative example of such a conversation is provided in Tab. 15. Each simulated interaction features a student working through a complex biology problem that requires multi-step reasoning and problem decomposition, with the tutor guiding the student through the process, making the dataset well-suited for our evaluation. Human annotators validated each simulated example to ensure factual accuracy and alignment with the socio-constructivist learning model.
The original authors used evaluation criteria that measured the accuracy of information (factual correctness), the relevance of the feedback to the current sub-problem and student errors (relevancy), the thoroughness of feedback in addressing all aspects of the sub-problem (completeness), and the impact of feedback on maintaining student interest (motivation).
In contrast, we concentrate solely on the non-affective components of the feedback, as prior studies have shown minimal influence of positive language and praise on student outcomes [53, 54]. Our primary objective is to develop evaluation criteria that assess the effectiveness of feedback in promoting conceptual understanding. Studies have consistently found that feedback elements which identify problems and errors and propose solutions are the most beneficial for student learning [55, 56, 57, 58]. Accordingly, our task is designed to evaluate an LLM’s ability to generate feedback that provides accurate problem diagnosis and rectification. While indirect feedback—such as hints or questions that avoid explicitly identifying the mistake or providing the answer—can foster meta-cognitive and conceptual learning, its effectiveness may diminish when students repeatedly struggle with the same sub-problem [59]. Accordingly, our evaluation also emphasizes the LLM’s ability to generate contextually appropriate feedback based on the nature of the student’s mistakes for a given sub-problem. In cases where the student makes repeated errors on the same sub-problem, the LLM should accurately recognize this pattern and respond with explicit error identification and corrective guidance. In this section, we explain how to get fine-grained information from the seed data collected by [51].
In the next subsection, we provide a brief overview of the relevant information contained in the seed data. In the following subsection, we describe how to extract relevant elements from this data to support evaluation based on our proposed criteria. Finally, we explain our procedure to sample a representative data covering diverse scenarios from this dataset.
Information Elements in Seed Data:
Consider a single response of the tutor from the example shown in Tab. 15. We focus on the second response from the tutor and provide fine-grained information associated with it in Tab. 13.
Table 13: T4EduPAE - A sample of the information elements annotated by the authors.
Tutor: Correct! A slower metabolic rate would help conserve energy. Now, let’s move on to the second subproblem: Identify a second derived feature that helps conserve energy in metabolism. Evaluation of the Student Response: Correct response from the student Subproblem: Identify a second derived feature that helps conserve energy in metabolism
The Evaluation of the Student Response field assesses the accuracy of the student’s most recent answer. Although the original seed data includes over eight categories for this entry, we focus on the following three: (a) Incorrect response, (b) Correct response, and (c) Partially correct response.
The Subproblem field indicates which sub-problem the feedback addresses. If the student’s response is correct, the feedback typically prompts them to proceed to a new sub-problem, and the Subproblem field reflects this next target. Conversely, if the response is incorrect or partially correct, the feedback remains focused on the current sub-problem, and the Subproblem entry corresponds to that same step.
Processing Data for Evaluating along our desired criteria:
To assess effectiveness in error identification and rectification, we extract these components from the feedback generated by the simulated tutor. This extraction is carried out using GPT-4o for automated annotation, with subsequent human verification to ensure the accuracy of the identified issues and corresponding corrective suggestions. Since the task does not require deep domain-specific knowledge, the manual verification was performed by one of the authors with strong proficiency in English. An illustrative example of the extracted elements from a single data point is provided in Tab. 14.
Table 14: T4EduPAE - A sample of the information elements extracted using GPT-4o.
Tutor: Insects are definitely affected due to the loss of native plants they specialize in. Additionally, think about the larger animals that rely on forest structure and plant diversity. For instance, how might birds and mammals be affected? Error Identification: The student response is not comprehensive in covering organisms such as birds and mammals that are also affected by the conversion. Error Rectification: Think about the larger animals that rely on forest structure and plant diversity. For instance, how might birds and mammals be affected?
To annotate whether a student has made repeated mistakes on a specific subproblem, we use the Evaluation of the Student Response and Subproblem fields to infer this state. Specifically, if the Subproblem value remains the same across two consecutive tutor responses and both responses evaluate the student’s answers as incorrect or partially correct, it indicates that the student is repeatedly struggling with the same subproblem.
Data Sampling: We sampled a total of $109$ conversation instances, ensuring that each context corresponds to a distinct problem. To study the impact of dialogue length, we selected approximately equal numbers of examples for conversation lengths in the set $\{5,7,\ldots,13\}$. Each length is an odd number, as conversations are structured to begin with a student query and end with a student response.
Furthermore, for most conversation lengths, we balanced the samples across the following five outcome categories: (a) ends with an incorrect response without a reattempt, (b) ends with an incorrect response after a reattempt, (c) ends with a correct response, (d) ends with a partially correct response without a reattempt, and (e) ends with a partially correct response after a reattempt.
B.4.4 Illustrative Example
An illustrative example is shown in Tab. 15. All samples for T4 are publicly available.
The sample input includes a multi-turn tutor-student dialogue in which the student is asked to identify features that help mammals survive in energy-scarce environments. The conversation is broken into subproblems to promote step-by-step reasoning. In this example, the student suggests “a larger body size” as an energy-conserving trait. The tutor must evaluate this response, determine its correctness, and provide pedagogically appropriate feedback that encourages deeper thinking without simply revealing the answer. The human reference demonstrates how to strike a balance between affirming partial correctness and guiding the student toward improved reasoning.
The model prompt is shown in Tab. 16.
Table 15: T4EduPAE - A sample of a tutor-student conversation and its corresponding pedagogical aligned response.
Tutor-student conversation (Sample input): Student: Q. Imagine you are a mammal species living in an environment where energy sources are scarce. Determine three derived features that could have arisen in response to the need for constant, high-level metabolism in such an environment. Tutor: Let’s break the problem into subproblems and tackle the subproblems one by one. Let’s begin with the first subproblem: Identify a derived feature that helps conserve energy in metabolism. Student: One feature could be a slower metabolic rate. Tutor: Correct! A slower metabolic rate would help conserve energy. Now, let’s move on to the second subproblem: Identify a second derived feature that helps conserve energy in metabolism. Student: Another feature could be a larger body size. Pedagogical aligned response (Human reference): Evaluation of the student response - Partially Correct Feedback - A larger body size could help conserve energy in some cases due to reduced surface area to volume ratio, which reduces heat loss. However, it depends on the specific environment and constraints. Can you think of another derived feature?
Table 16: T4EduPAE – Model prompt.
You are an tutoring agent, an AI-powered expert chatbot designed to help assist teachers in constructing the right pedagogical response for a conversation thread between the teacher and the student. In a particular conversation thread, the student asks the teacher a question. The teacher adopts the strategy of dissociating the question into several sub-questions, and then tackling each sub-question one by one. The teacher does this by asking the student a sub-question, and then helping the student to solve it until the student is able to answer it correctly. The teacher then moves on to the next sub-question, and so on, until all sub-questions are solved. In a given conversation thread, the teacher may have already helped the student solving several sub-questions. You are required to assess the latest sub-question being tackled by the teacher and its associated student response(s), and provide a pedagogically aligned response to the teacher. You are guaranteed that the conversation thread provided to you always ends with a student response. This pedagogically aligned response must have the following elements: 1. Evaluation of the student response: Evaluate the student response into one of the following categories: - a) Incorrect response: The student response is incorrect. - b) Correct response: The student response is correct and answers the sub-question. - c) Partially correct response: Either the student response is partially correct, or the student response is correct but does not completely answer the sub-question. 2. Repeated mistake: If the student response is incorrect or partially correct, check if the student has made a mistake for the same sub-question before. If the student has made a mistake for the same sub-question before, answer "Yes". Otherwise "No". 3. Feedback: Provide a holistic feedback to the student in a plain text format. The feedback must identify mistakes if the student response is incorrect or partially correct. Furthermore, identify the mistake indirectly without explicitly pointing it out if the student makes an error for the first time for the corresponding sub-problem. Otherwise, highlight the mistake explicitly. The feedback must also provide a hint when the student response is incorrect or partially correct. Furthermore, the hint must indirect when a mistake has been made for the first time for a sub-problem. Otherwise, provide an explicit hint that would help student rectify their mistakes directly. Include a praise only when the student performs the task correctly or partially correctly. Include motivation and encouragement when the student is making multiple mistakes for the sub-problem being targeted. Finally, if the student response is correct, suggest the next sub-question if applicable. The pedagogically aligned response must be provided in the following format: Evaluation of the student response: <evaluation> Repeated mistake: <Yes/No> Feedback: <Holistic Feedback provided as a plain text whose content may include mistake identification (if required), mistake rectification (if required), praise (if required), encouragement and motivation (if required), and sub-question transition (if required).>
B.4.5 Evaluation Rubric
The detailed evaluation rubric is provided below:
  1. 1.
    Evaluation of the student response: Evaluating the response generated by the LLM in accurately identifying whether the student response is correct, incorrect or partially correct.
  2. 2.
    Repeated mistake: Whether the LLM generated response accurately identifies if the student’s reattempt along a problem in incorrect / partially correct.
  3. 3.
    Error Identification: Whether the error identified by the LLM generated response is consistent with that of the reference if present. For cases where student response is correct, the LLM must not generate “N/A”.
  4. 4.
    Error Rectification: Whether the elements that describe the rectification of the error is consistent with that of the reference if present. For cases where student response is correct, the LLM must not generate “N/A”.
  5. 5.
    Error Identification Type: Whether the feedback involves direct error identification when the student makes multiple inaccurate / partially accurate attempts along a sub-problem and involve indirect problem identification otherwise
  6. 6.
    Error Rectification Type: Whether the feedback involves direct rectification of the errors when the student makes multiple inaccurate / partially accurate attempts along a sub-problem and involve indirect problem rectification otherwise
  7. 7.
    Sub-question Transition: Whether the next sub-question in the LLM generated response is consistent with that of the reference. If all the sub-problems are solved or if the student makes an error along current one, then the feedback must not involve this aspect.
B.4.6 Checklist-mapped Reference
A prompt for generating the checklist-mapped reference was unnecessary, as it was produced directly from the model output.
Instead, to evaluate the quality of the generated output according to the rubric , we extracted comparable information from the reference tutor data. Section B.4.3 already describes how to identify elements that describe a problem and corresponding solutions. Additionally, it provides a method for determining whether the student is struggling with the current sub-problem—specifically by checking for repeated errors across attempts. This Repeated mistake attribute is used to assign appropriate values to the Error Identification type and Error Rectification type. These types are set to Indirect when the student is attempting the sub-problem for the first time. If the Repeated mistake flag is active, both types are instead set to Direct. Notably, these properties are only assigned when the student’s response contains an error. For the Sub-question Transition property, we directly use the value provided in the original dataset, which is only set when the student gives a correct answer but the overall problem remains unsolved. An example of the checklist-mapped human reference from Tab. 15 is shown in Tab. 17.
Table 17: T4EduPAE - Checklist-mapped reference.
1. Evaluation of the student response: Partially correct response 2. Repeated mistake: No 3. Error Identification: The student response does not fully address the suitability of a larger body size as a derived feature for conserving energy in all environments. 4. Error Rectification: Can you think of another derived feature? 5. Error Identification Type: Indirect 6. Error Rectification Type: Indirect 7. Sub-question Transition: N/A
B.5 T5EduFG
B.5.1 Task Definition
The objective of this task is to evaluate whether an LLM can assess student responses with the same discernment as a trained instructor or expert. In this setup, we consider an assignment designed by the course instructor, whose response takes the form of a long-form essay. Although the instructor provides guidelines that students are expected to follow, these are intentionally not exhaustive—ensuring that students are not simply spoon-fed a detailed checklist of requirements. Instead, students are expected to adhere to additional implicit expectations informed by the course material covered thus far. While instructors are skilled at evaluating responses based on these unstated criteria, this task aims to determine whether an LLM can perform such evaluations effectively.
More concretely, the assignment includes both public and private requirements. While the instructor is aware of both, the LLM is given only the public requirements and the course syllabus, and must use these to also evaluate the response against the private requirements.
B.5.2 Task Significance
Our motivation to assess LLMs’ ability to evaluate student responses beyond student-facing rubric stems from the practical insight that instructors often do not rely solely on a rubric Instead, they apply holistic judgments that incorporate broader, often implicit, expectations [60, 61]. This is particularly evident in writing assignments, where public rubrics tend to be too general to support nuanced evaluation. Instructors frequently augment them with private, instance-specific criteria to enable more accurate and context-sensitive scoring [62]. Building on these observations, we seek to determine whether LLMs can similarly move beyond surface-level guidelines to produce evaluations that align with instructor judgment.
To study this, we use responses to an assignment, where the student-facing (public) rubric is subsumed by the private rubric used by the instructor. This setup allows us to test the LLM’s capacity to evaluate in a manner consistent with expert grading practices, including criteria not explicitly stated.
B.5.3 Data Acquisition
Data Description:
For this task, we use data from the Economics 101 course at the University of Michigan, Ann Arbor [63]. The assignment presents a scenario in which “an increase in the minimum wage in San Francisco could lead to increased adoption of automation.” To counter this, two policy options are proposed: (a) a tax on automation and (b) a ban on automation. Students are asked to write a persuasive letter outlining the economic implications of the wage increase and to argue against one of the proposed policies using concepts and tools covered in the course. The student-facing rubric is shown in Tab. 18 - it is clearly evident that the student-facing rubric is high-level and cannot be used as a criteria for robustly assigning the scores.
Table 18: T5EduFG - Public/student-facing rubrics for the ECON 101 Assignment.
Understanding Check whether all the relevant economic concepts central to the policies and markets are identified and correctly defined in a way that exceeds expectations for the course. Identify the missing concepts and concepts that are incorrectly defined. Building upon their definitions, assess whether the writer correctly connects the relevant concepts and markets to one another demonstrating an understanding that is sophisticated for the course. Identify the missing connections and connections that are incorrectly made. Critical Thinking Assess whether the writer accurately interprets and articulates the economics within the source in a sophisticated manner while predominantly summarizing the source. Identify the missing interpretations and interpretations that are incorrectly made. Identify if the essay lacks citations or has incorrect citations. Assess whether the author provides insightful articulation of the issues facing one of the proposed solution. Check whether all the market interactions are explored coming to the solution indicating that the proposed solution is not economically sound. Determine whether the author accurately interweaves each economic concept present in the proposal into their articulation of the downsides. Specify the missing interactions and interactions that are incorrectly made. Determines the concepts that are missing in the articulation of the downsides or incorrectly defined. Response Alignment with Audience Assess whether the explanation aligns with the recommended audience. Check whether the recommendations are inconsistent with the target audience - for instance, recommending government action when the audience is a producer. Determine whether the explanations are too advanced or too simple for the specific audience.
In contrast, the instructor also uses a set of private evaluation rubrics to enable more fine-grained and robust assessment of the essays. Examples include:
  • Identifying the shift from a non-binding to a binding price floor
  • Recognizing key concepts such as price floor/minimum wage, binding vs. non-binding constraints, substitutes, and labor vs. automation
  • Explaining how an increase in the minimum wage could reduce the supply of final goods/services due to higher input costs or firm closures
In total, there are $20$ such private rubric items used to guide a more nuanced evaluation of the essay content.
Each student in the course submits an essay-based assignment, which is then evaluated by the instructor or teaching assistant using the private set of rubrics to guide their feedback. Although the feedback may be unstructured, it can be parsed into a structured format by mapping it to Yes/No responses for each item in the private rubric list.
Feedback Parsing:
For a given unstructured feedback, we use GPT-4o to extract information along each rubric item as shown before. For instance, the prompt to extract above items would be:
  • Does the feedback indicate that the following statement is missing in the essay: Identification of the change from nonbinding to binding price floor?
  • Does the feedback identifies that the following concepts are missing from the essay: Price floor/min wage, binding/non-binding, substitutes, labor/automation
  • Does the feedback indicate that the essay contains an error along the following aspect: The increase of minimum wage will lead to a decrease in supply for final goods/services due to an increase in input prices or firms going out of business?
An affirmative response to a prompt implies that the corresponding rubric item is not properly addressed in the essay. This forms the checklist-mapped reference for each feedback.
Sampling:
To sample 100 data points—including student responses, instructor feedback, and the corresponding checklist-mapped references—we selected those with the longest instructor feedback, based on word count. This is because essays with longer instructor feedback typically contain more errors, requiring more careful analysis to identify and address them.
B.5.4 Illustrative Example
Due to the proprietary nature of the data, we cannot provide a specific example. However, an example of instructor feedback is provided in Tab. 19 for a student written essay and the model prompt is shown in Tab. 20:
Table 19: T5EduFG - An example of instructor feedback.
Did not define price floor or distinguish binding from nonbinding. Need to state that automation & labor are substitutes in consumption. Need to be more clear in your explanation of the effects on the three markets (what happens in the final goods and service market as a result of the binding minimum wage?) Additionally, review the difference between quantity supplied/demanded and supply/demand. You state that businesses might reduce their demand for labor following a binding minimum wage, however, when the minimum wage becomes binding, firm’s quantity demanded decreases, while the quantity supplied of labor will increase. Then, since automation and labor are substitutes in consumption, the demand for the automation will increase. This increases the price for automation. As a result, the supply of goods and services actually decreases - input prices for labor (manual or automated) has increased No in-text citations or references (sources) at the end of letter. Additionally, instead of quoting directly, you should paraphrase and devote more of your word count to the economic analysis While you successfully acknowledged that a ban or tax creates deadweight loss/ reduces total surplus, it was essential that you focused on one of the policies, and explain the impact on all THREE markets (using supply and demand analysis) We are assuming the minimum wage already increased, it is not a hypothetical scenario. Additionally, need to focus on opposing one of the two policies, you repeatedly mentioned both, just need to focus on one.
Table 20: T5EduFG – Model prompt.
Given the essay prompt and the concepts covered in the course so far, please provide a detailed and comprehensive feedback along the following aspects: Understanding Check whether all the relevant economic concepts central to the policies and markets are identified and correctly defined in a way that exceeds expectations for the course. Identify the missing concepts and concepts that are incorrectly defined. Building upon their definitions, assess whether the writer correctly connects the relevant concepts and markets to one another, demonstrating an understanding that is sophisticated for the course. Identify the missing connections and connections that are incorrectly made. Critical Thinking Assess whether the writer accurately interprets and articulates the economics within the source in a sophisticated manner while predominantly summarizing the source. Identify the missing interpretations and interpretations that are incorrectly made. Identify if the essay lacks citations or has incorrect citations. Assess whether the author provides insightful articulation of the issues facing one of the proposed solutions. Check whether all the market interactions are explored in coming to the solution, indicating that the proposed solution is not economically sound. Determine whether the author accurately interweaves each economic concept present in the proposal into their articulation of the downsides. Specify the missing interactions and interactions that are incorrectly made. Determine the concepts that are missing in the articulation of the downsides or incorrectly defined. Response Alignment with Audience Assess whether the explanation aligns with the recommended audience. Check whether the recommendations are inconsistent with the target audience—for instance, recommending government action when the audience is a producer. Determine whether the explanations are too advanced or too simple for the specific audience.
B.5.5 Evaluation Rubric
In this section, we present the rubric used for evaluation. Each rubric item is annotated with the type of issue it addresses. Specifically, (missing) indicates whether the feedback correctly identifies the absence of a corresponding statement, while (error) assesses whether the feedback accurately detects an error in the corresponding element. To measure whether a certain mistake has been accurately detected, we check whether the checklist-mapped response entry matches with the reference one.
  1. 1.
    Identifies the change from nonbinding to binding price floor (missing): Does the feedback accurately indicate that the following statement is missing in the essay: Identification of the change from nonbinding to binding price floor?
  2. 2.
    Identifies the concepts of price floor/min wage, binding/non-binding, substitutes, labor/automation (missing): Does the feedback accurately identify that the following concepts are missing from the essay: Price floor/min wage, binding/non-binding, substitutes, labor/automation?
  3. 3.
    Mentions that the increase of minimum wage causes a decrease in the demand for automation (error): Does the feedback accurately mention that the essay contains erroneous information such as: the increase of minimum wage causes a decrease in the demand for automation?
  4. 4.
    Defines the concepts of price floor/min wage, binding/non-binding, substitutes, labor/automation (missing): Does the feedback accurately identify that the definitions of the following concepts are missing from the essay: Price floor/min wage, binding/non-binding, substitutes, labor/automation?
  5. 5.
    (Only when the student is writing against the ban) States that a ban is essentially a [quota of zero] (missing): Does the feedback accurately indicate that the following statement is missing in the essay: A ban is essentially a [quota of zero]?
  6. 6.
    Specify that minimum wage is a price floor (missing): Does the feedback accurately indicate that the following statement is missing in the essay: Minimum wage is a price floor?
  7. 7.
    Treating the labor market/final good market as a homogeneous entity without distinguishing different workers or labor-made/automation-made goods (error): Does the feedback accurately identify that different workers or labor-made/automation-made goods are distinguished, making the labor market/final good market a non-homogeneous entity?
  8. 8.
    State that the firm has lower costs of production after the minimum wage increase (missing): Does the feedback accurately indicate that the following statement is missing in the essay: The firm has lower costs of production after the minimum wage increase?
  9. 9.
    Mentions that automation and labor are substitutes in consumption (missing): Does the feedback accurately indicate that the following statement is missing in the essay: Automation and labor are substitutes in consumption?
  10. 10.
    Explains that the increase of minimum wage will lead to a decrease in supply for final goods/services due to an increase in input prices or firms going out of business (error): Does the feedback accurately indicate that the essay contains an error along the following aspect: The increase of minimum wage will lead to a decrease in supply for final goods/services due to an increase in input prices or firms going out of business?
  11. 11.
    Explains that the increase of minimum wage will lead to a decrease in supply for final goods/services due to an increase in input prices or firms going out of business (missing): Does the feedback accurately indicate that the essay lacks an explanation along the following aspect: The increase of minimum wage will lead to a decrease in supply for final goods/services due to an increase in input prices or firms going out of business?
  12. 12.
    State that with tax or ban on automation, the demand for labor increases (missing): Does the feedback accurately indicate that the following statement is missing in the essay: Tax or ban on automation results in increased demand for labor?
  13. 13.
    State that with tax or ban on automation, the supply of the final goods/service decrease (missing): Does the feedback accurately indicate that the following statement is missing in the essay: Tax or ban on automation results in decreased supply of the final goods/service?
  14. 14.
    State that with the binding minimum wage, the quantity demanded for the labor will decrease (missing): Does the feedback accurately indicate that the following statement is missing in the essay: The binding minimum wage decreases the quantity demanded for the labor?
  15. 15.
    Provide reason for the shift when identifies a curve shift or change in demand/supply (missing): Does the feedback accurately indicate that the reason for the shift is missing when identifying a curve shift or change in demand/supply?
  16. 16.
    Explain conceptually instead of solely rely on shifting the supply and demand curve (missing): Does the feedback accurately indicate that the essay lacks conceptual explanation and solely relies on shifting the supply and demand curve?
  17. 17.
    Identify the change in quantity supplied/demanded instead of supply/demand (missing): Does the feedback accurately indicate that the essay does not identify the change in quantity supplied/demanded instead of supply/demand?
  18. 18.
    Propose a solution on their own to the problem of the increasing minimum wage (missing): Does the feedback accurately indicate that the solution to the problem of increasing minimum wage is not proposed in the essay?
  19. 19.
    Does not explain concepts the readers are expected to know from the prompt (i.e. demand, supply, consumer surplus, producer surplus, and efficiency). (error): Does the feedback accurately indicate that the essay explains concepts already known to the readers such as demand, supply, consumer surplus, producer surplus, and efficiency?
  20. 20.
    Address all the three markets and the inter-market effects at play (missing): Does the feedback accurately identify that the essay fails to address all the three markets and the inter-market effects at play?
For each item, “Yes” indicates that the essay has a problem along the corresponding element as indicated by the feedback. For each response generated by the LLM, we parse it into a similar format and measure the extent of alignment between the answers along each checklist item.
B.5.6 Checklist-mapped Reference
A prompt for creating the checklist-mapped reference is not necessary in this task, as all model responses are binary (“Yes” or “No”) and easily verifiable. An example of the checklist-mapped reference of the instructor feedback in Tab. 19 is presented in Tab. 21.
Table 21: T5EduFG - Checklist-mapped reference.
1. Identifies the change from [nonbinding] to [binding] [price floor] (missing): No 2. Identifies the concepts of price floor/min wage, binding/non-binding, substitutes, labor/automation (missing): Yes 3. Mentions that the increase of minimum wage causes a decrease in the demand for automation (error): No 4. Defines the concepts of price floor/min wage, binding/non-binding, substitutes, labor/automation (missing): Yes 5. (Only when the student is writing against the ban) States that a ban is essentially a [quota of zero] (missing): No 6. Specify that minimum wage is a price floor (missing): No 7. Treating the labor market/final good market as a homogeneous entity without distinguishing different workers or labor-made/automation-made goods (error): No 8. State that the firm has lower costs of production after the minimum wage increase (missing): No 9. Mentions that automation and labor are substitutes in consumption (missing): Yes 10. Explains that the increase of minimum wage will lead to a decrease in supply for final goods/services due to an increase in input prices or firms going out of business (error): Yes 11.
B.6 T6HealthCNG
B.6.1 Task Definition
Clinical note generation is the task of producing well-structured, accurate, and comprehensive clinical notes based on patient-doctor dialogue during a clinical encounter [64]. In this task, we focus on generating SOAP notes, as they are the most common way to document medical interactions [65]. SOAP notes consist of four key sections: Subjective (patient-reported symptoms), Objective (clinician-observed data), Assessment (diagnoses or impressions), and Plan (treatment or follow-up steps). The objective is to generate accurate, structured, and useful SOAP notes from unstructured patient and doctor conversations. Automating this process would greatly reduce the clinician workload and enhance the quality of medical documentation. In using LLMs to perform this task, we input a the transcript of a conversation between a healthcare professional and patient. The output is a SOAP note that includes the key information from the conversation. An example note is shown in Tab. 22.
B.6.2 Task Significance
Clinical documentation plays an important role in effective, high-quality care by supporting accurate diagnoses and standardizing communication among healthcare providers. However, the generation of structured clinical notes is time-consuming and prone to human error. Studies show that physicians can spend over an hour documenting a single clinical visit and nearly two hours on electronic health record (EHR) tasks for every hour of direct patient care [66]. Given the demanding schedules of healthcare professionals, this burden increases the risk of misremembered or inaccurately entered information. Automating this process with large language models (LLMs) can improve consistency and streamline workflows, ultimately reducing patient wait times and enhancing the quality of care.
B.6.3 Data Acquisition
We use the ACI-Bench dataset, a benchmark for clinical note generation with 207 annotated clinical encounters. Each sample includes a transcript of a patient-doctor interaction and its corresponding SOAP structured note written by the doctor [67]. Transcripts were created using three methods: (1) a virtual assistant, in which the doctor had to use explicit terms (e.g. “Hey Dragon show me the diabetes lab”) during the visit; (2) a virtual scribe, automated or otherwise, which assisted in note creation without distracting in-person doctor-patient interactions; and (3) an ambient clinical intelligence (ACI), which created the transcript without interrupting natural conversation flow between the patient and physician.
From these 207 samples, we selected a final subset of 100 on difficulty and diversity consistent with the rationale in §3.2. Difficulty was measured using the length of the human reference. Longer SOAP notes will contain more information about the patient, treatment, and analysis, making diagnosis a more difficult task. Additionally, ACI-Bench was originally collected in a diverse manner. Not only were transcripts created in a variety of methods, the content of the SOAP note itself was also diverse. For example, some notes contained Chief Complaints, but others did not; other notes had longer History of Present Illness sections, while others included symptom context in the Assessment or Plan sections.
B.6.4 Illustrative Example
An illustrative example is shown in Tab. 22. All samples for T6 are publicly available.
The sample input consists of a multi-turn conversation capturing the doctor’s history-taking process and the patient’s responses about symptoms, routines, and relevant background. In this example, the patient is a teenage girl accompanied by her mother—presents for an acne evaluation. The human reference is a structured clinical note in SOAP format. The note reflects standard clinical reasoning, summarizes key details, and serves as a reference for future clinical visits.
The model prompt shown in Tab. 23 was taken directly from the ACI-Bench benchmark [67]. Although the prompt does not explicitly specify the four SOAP sections, the paper explains that it was strongly adapted from the SOAP note format: the History of Present Illness aligns with the Subjective section, the Exam and Results correspond to the Objective section, and the Assessment and Plan are combined into the final section. This structure also justifies our use of checklist items that incorporate SOAP elements in Section B.6.5.
Table 22: T6HealthCNG - A sample of a transcript of a conversation between a patient and doctor and its corresponding SOAP note.
Transcript (Sample input): [doctor] kayla ward , date of birth , 4/28/07 . mrn 3-8-4-9-2-0 . she’s here for a new visit with her mother for acne located on the face , which started about two years ago and is present most every day . she has been using persa-gel and washing regularly , which is somewhat helpful . there are no associated symptoms including itching , bleeding , or pain . no additional past medical history . she lives with her parents and sister . they have a dog , bird , and bunnies . she is in 7th grade . she plays basketball and volleyball and tap . she wears sunscreen in the summer , spf 30 . no additional family history . hi kayla , i’m dr. juan price . i hear you are starting to get some acne on the face . how about the chest and back ? [patient] it’s not too bad . [doctor] so , it’s not bad on the chest or back . you’ve used some over the counter items like washes and persa-gel ? [patient] yeah . [doctor] do those seem to be helping ? [patient] yes , i think so , a little bit . [doctor] good . what’s your skin care routine like now ? [patient] do you wan na know , like , the things i currently use ? [doctor] yes . what do you do for your acne in the morning ? and then what do you do at nighttime ? [patient] i wash my face , more like i wipe it down in the morning . then at night i use an elf facial cleanser called the super clarity cleanser . i finish with a toner and then the persa-gel . SOAP note (Human reference): CHIEF COMPLAINT New acne evaluation. HISTORY OF PRESENT ILLNESS Kayla Ward is a 15-year-old female who presents for new patient evaluation of acne located on the face. She is accompanied by her mother today. Kayla states her acne started approximately 2 years ago and it is present almost every day. The patient’s mother notes that the most significant acne flares started in the fall when she was playing school sports. It does not tend to flare with her periods. Kayla reports that today is a good day for her acne. She denies any significant acne present on the chest or back. There are no associated symptoms, including no itching, bleeding, or pain. The patient has been washing her face regularly. Her acne regimen includes washing her face in the morning with Persa-Gel and at night e.l.f. SuperClarify Cleanser along with toner and Persa-Gel. This regimen is somewhat helpful. She wears sunscreen in the summer SPF 30. …
Table 23: T6HealthCNG - Model prompt.
Summarize the conversation to generate a clinical note with four sections: HISTORY OF PRESENT ILLNESS, PHYSICAL EXAM, RESULTS, ASSESSMENT AND PLAN. The conversation is:
B.6.5 Evaluation Rubric
In T6, we did not consult domain experts as in previous tasks, but created our rubric independently through extensive research. Our rubric items and definitions were verified across multiple reputable online sources [64, 65] and are still expert-level. The rubric comprises 29 checklist items that captures the key information present in a SOAP note. The detailed checklist items are listed as follows:
  • Subjective (S) Section
    1. 1.
      Reason for Patient Office Visit or Hospitalization: The primary reason for the patient’s visit. Identifies the most pressing issue if multiple complaints are present. Uses concise and medically appropriate language.
    2. 2.
      Patient’s Age
    3. 3.
      Patient’s Sex
    4. 4.
      Patient’s Reason for the Visit
    5. 5.
      Onset (if applicable): When the complaint started.
    6. 6.
      Location: The exact location the Chief Complaint happened.
    7. 7.
      Duration: How long the complaint has persisted.
    8. 8.
      Character: How the patient describes the Chief Complaint.
    9. 9.
      Alleviating & Aggravating Factors: What makes the issue better or worse.
    10. 10.
      Radiation: Whether symptoms move or stay in one spot.
    11. 11.
      Temporal Factor: Whether or not the Chief Complaint is worse (or better) at a certain time of the day.
    12. 12.
      Severity: The rating from the patient about the Chief Complaint using a scale of 1 to 10, 1 being the least pain, 10 being the worst pain.
    13. 13.
      Relevant Medical History: Any relevant medical history, including past diagnoses, surgeries, and hospitalizations.
    14. 14.
      Surgical History (if applicable)
    15. 15.
      Family History
    16. 16.
      Home and Environment: The patient’s living situation, relationships with family/roommates, and sense of safety/stability at home.
    17. 17.
      Education: The patient’s current level of schooling, academic performance, or school engagement.
    18. 18.
      Employment: The patient’s job status, job satisfaction, work hours, or financial independence.
    19. 19.
      Eating Habits: The patient’s diet quality, body image concerns, or disordered eating patterns.
    20. 20.
      Activities: The patient’s hobbies, friends, online/social media use, or after-school/work activities.
    21. 21.
      Drugs: Use of alcohol, tobacco, marijuana, or other substances.
    22. 22.
      Sexuality: The patient’s sexual activity, orientation, or gender identity.
    23. 23.
      Suicide/Depression: The patient’s mood, self-harm, suicidal ideation, or prior mental health diagnoses or treatments.
    24. 24.
      Review of Systems (ROS): Describes an inventory of the body systems to identify signs and/or symptoms which the patient may be experiencing. The body systems must fall into one of the 14 systems: Constitutional symptoms (i.e. fever, weight loss, vital signs); Eyes; Ears, nose, mouth, throat; Cardiovascular; Respiratory; Gastrointestinal; Genitourinary; Musculoskeletal; Integumentary; Neurological; Psychiatric; Endocrine; Hematologic/Lymphatic; and Allergic/Immunologic. Document both positive and pertinent negatives for each system reviewed.
  • Objective (O) Section
    1. 25.
      Vital Signs: BP, HR, RR, Temp, SpO2, weight, height (if relevant).
    2. 26.
      Physical Examination Findings: The basic systems of cardiac and respiratory, affected systems, possible involvement of other systems, pertinent normal findings and abnormalities.
    3. 27.
      Other Objective Data (if applicable): Results from laboratory and other diagnostic tests already completed.
  • Assessment (A) Section
    1. 28.
      Diagnosis & Clinical Impression: Provides a tentative diagnosis, assessment of patients status based on subjective and objective findings, a list of other possible diagnoses usually in order of most likely to least likely. The assessment will also include possible and likely etiologies of the patient’s problem. It is the patient’s progress since the last visit, and overall progress towards the patient’s goal from the physician’s perspective.
  • Plan (P) Section
    1. 29.
      Treatment & Management Plan: States what the health care provider will do to treat the patient’s concerns—such as ordering further labs, radiological work up, referrals given, procedures performed, medications given and education provided. The plan will also include goals of therapy and patient-specific drug and disease-state monitoring parameters. This should address each item of the differential diagnosis. For patients who have multiple health problems that are addressed in the SOAP note, a plan is developed for each problem and is numbered accordingly based on severity and urgency for therapy. A note of what was discussed or advised with the patient as well as timings for further review or follow-up are generally included. This part is often grouped together with Assessment.
B.6.6 Checklist-mapped Reference
The checklist-mapped references for both the human reference and model output were constructed with a similar approach to that used in T1 (see Section B.1.6). However, in T6, we grouped 29 items into five groups based on (1) the expected length of the model’s response and (2) the SOAP category the item belonged to (e.g., Subjective, Objective, etc.). This is to help the model focus on distinct clinical sections and break down complex notes into meaningful, manageable parts. The prompt for extracting an individual checklist item is presented in Tab. 24. A checklist-mapped reference of the sample in Tab. 22 is shown in Tab. 25.
To assess the quality of the checklist-mapped reference, we selected an additional 30 difficult and diverse samples, different from the main dataset, and conducted human verification to examine the faithfulness of the model response. Details of this experiment can be found in LABEL:ap:validation-of-checklist.
Table 24: T6HealthCNG - Prompt for extracting checklist-mapped reference.
You are an experienced doctor reviewing clinical notes to identify key medical information. Given a clinical note, extract the Reason for Patient Office Visit or Hospitalization information. Extract crucial related information as completely as possible. The Reason for Patient Office Visit or Hospitalization Clearly states the primary reason for the patient’s visit. If multiple complaints are present, it identifies the most pressing issue. Example Format: [Reason for Patient Office Visit or Hospitalization]. If no related information is mentioned in the clinical note, state “N/A”. If no related information is mentioned in the clinical note, state “N/A”. This is the clinical note:
Table 25: T6HealthCNG - Checklist-mapped reference.
1. Reason for Patient Office Visit or Hospitalization: New acne evaluation 2. Patient Age: 15 3. Patient Sex: female 4. Patient Reason for the Visit: New acne evaluation 5. Onset: Approximately 2 years ago 6. Location: Face, primarily on the forehead, with also some on the central cheeks and chin 7. Duration: 2 years 8. Character: Present almost every day, primarily on the forehead, central cheeks, and chin 9. Alleviating and Aggravating Factors: Flares in the fall during school sports, does not flare with periods, regimen is somewhat helpful 10.
B.7 T7ChemMDG
B.7.1 Task Definition
Molecule description generation involves creating accurate and structured natural language descriptions of molecular structures based on their SMILES (Simplified Molecular Input Line Entry System) representations [68]. SMILES encodes molecules as linear strings that represent atoms, bonds, rings, and branching patterns, and serves as a textual representation of molecular graphs. The objective is to translate these symbolic sequences into natural language descriptions that capture key structural and chemical features of the molecule. When using LLMs to complete this task, the input consists of a SMILES string (e.g., “CC(=O)OC1=CC=CC=C1C(=O)O”). The output is a molecule description a professional might read. An example is shown in Tab. 26.
B.7.2 Task Significance
Chemical databases contain tens of millions of molecules, each represented by complex SMILES strings that are not easily interpretable by humans. Experts often find it time-consuming and challenging to infer the functional class or properties of a molecule solely from its SMILES representation, especially for more complex structures. Automating the translation of SMILES into structured, ontological, and natural language descriptions can significantly enhance the accessibility and usability of chemical data. This advancement supports key fields such as drug discovery, materials science, and chemical education by enabling quicker understanding and communication of molecular information [69].
For example, models such as MolT5 have been specifically developed to complete this task. MolT5 leverages transformer-based architectures to improve the quality and informativeness of molecular descriptions, facilitating a deeper understanding of molecular structures and their properties [70]. Such models not only accelerate the discovery phase but also improve collaboration across multidisciplinary research teams.
B.7.3 Data Acquisition and Preprocessing
We use the ChEBI-20 dataset originally collected for the Text2Mol task [69]. This dataset was created by collecting compound annotations from the ChEBI database [^20] https://www.ebi.ac.uk/chebi/, which were scraped from PubChem [^21] https://pubchem.ncbi.nlm.nih.gov/. Descriptions shorter than 20 words were excluded to ensure sufficient detail. The resulting dataset contains 33,010 pairs of SMILES strings and their corresponding textual descriptions. Both ChEBI and PubChem are specialized, domain-specific chemical resources commonly utilized by chemists, aligning well with the expert-level knowledge demanded by this task.
To specifically evaluate the capability of LLMs in generating detailed molecule descriptions, we selected a small subset from ChEBI-20 based on difficulty. In T7, difficulty is approximated by the length of the human-written reference, with longer descriptions generally indicating more complex or detailed explanations. We selected 100 samples whose reference descriptions exceed 500 characters to focus on more challenging examples.
B.7.4 Illustrative Example
An illustrative example is shown in Tab. 26. All samples for T7 are publicly available.
The sample input consists of a SMILES string representing the chemical structure. The human reference is a detailed description that captures both the molecular structure and pharmacological function of the compound. It uses precise chemical nomenclature and is clear and concise.
The model prompt is presented in Tab. 27. This prompt was carefully designed in collaboration with domain experts to ensure both generality and relevance across a broad range of chemical compounds. To align with the ChEBI dataset, which stands for “Chemical Entities of Biological Interest,” we incorporated the phrase “Chemicals of Biological significance” into the prompt. However, we deliberately avoided including the acronym “ChEBI” itself to prevent the model from relying on any dataset-specific shortcuts.
Table 26: T7ChemMDG - A sample of a SMILES string and its corresponding molecule description.
SMILES string (Sample input): “C1CN2C(=CC=C2C(=O)C3=CC=CC=C3)C1C(=O)O” Molecule description (Human reference): The molecule is a racemate comprising equimolar amounts of (R)-(+)- and (S)-(-)-5-benzoyl-2,3-dihydro-1H-pyrrolizine-1-carboxylic acid. While only the (S)-(-) enantiomer is a COX1 and COX2 inhibitor, the (R)-(+) enantiomer exhibits potent analgesic activity. A non-steroidal anti-inflammatory drug, ketorolac is mainly used (generally as the tromethamine salt) for its potent analgesic properties in the short-term management of post-operative pain, and in eye drops to relieve the ocular itching associated with seasonal allergic conjunctivitis. It was withdrawn from the market in many countries in 1993 following association with haemorrhage and renal failure. It has a role as a cyclooxygenase 2 inhibitor, a cyclooxygenase 1 inhibitor, a non-steroidal anti-inflammatory drug and an analgesic. It contains a (R)-ketorolac and a (S)-ketorolac. It is a conjugate acid of a ketorolac(1-).
Table 27: T7ChemMDG – Model prompt.
You are a chemical researcher in charge of writing descriptions of Chemicals of Biological significance given their Simplified Molecular Input Line Entry System (SMILES) structure. Use domain specific terminology and.specific molecule names. Be as specific as possible. Please provide your description in paragraph format. Here is the SMILES structure:
B.7.5 Evaluation Rubric
Furthermore, we develop a fine-grained, checklist-based evaluation rubric in collaboration with a chemistry domain expert to ensure comprehensive coverage of key information in a molecule description. The expert not only advised us on the selection of rubric items, but also annotated a subset of human-written reference descriptions to identify which specific elements were explicitly or implicitly mentioned. These annotations informed the basis for the design of our checklist and guided the formulation of model prompts, ensuring that generated descriptions capture the same level of completeness and domain accuracy as the human references.
The final rubric comprises of six checklist items that capture key information typically found in professional molecule descriptions. The last sentence of each definition (starting with “Usually, but not always,”) is included to support the construction of the checklist-mapped reference described in Section B.7.6. The detailed checklist items are listed as follows:
  1. 1.
    Structure: The chemical composition of the molecule. Usually, but not always, this description begins with “The molecule is”.
  2. 2.
    Biological Function and Applications: The molecule’s functions in living organisms, including its interactions, effects on biological processes, and applications in the pharmacological domain. Usually, but not always, this description begins with “It has a role as”.
  3. 3.
    Chemical compound classifications: The specific group(s) of atoms that are used to categorize chemical compounds based on their structural characteristics, functional groups, or biological origin. Usually, but not always, this description begins with “It is a member of”.
  4. 4.
    Conjugate base (if applicable): The species that remains after an acid donates a proton (H+) in a chemical reaction. Usually, but not always, this description begins with “It is a conjugate base of”.
  5. 5.
    Conjugate acid (if applicable): The species that remains after an base accepts a proton (H+) in a chemical reaction. Usually, but not always, this description begins with “It is a conjugate acid of”.
  6. 6.
    Origin (if applicable): The molecule or chemical compound that has been extracted or obtained from a specific natural resource. Usually, but not always, this description begins with “It is isolated from”.
B.7.6 Checklist-mapped Reference
In line with the method described in T1, we generated checklist-mapped references for the human reference and model outputs with our evaluation rubric. However, with only six checklist items, we were able to create one group with one model extraction prompt. The prompt for extracting an individual checklist item is presented in Tab. 28 and the model output extraction prompt is available in our public GitHub repository. The checklist-mapped reference for the sample in Tab. 26 is shown in Tab. 29.
Table 28: T7ChemMDG - Prompt for extracting checklist-mapped reference.
You are assisting a chemical researcher in extracting key information from a molecule description. Given a molecule description, extract the Structure: the chemical composition of the molecule. Extract crucial related information as completely as possible. Usually, but not always, this description begins with “The molecule is”. Only respond with extracted text from the description related to the structure. If the structure is not mentioned, state “N/A”. This is the molecule description:
Table 29: T7ChemMDG - Checklist-mapped reference.
1. Structure: a racemate comprising equimolar amounts of (R)-(+)- and (S)-(-)-5-benzoyl-2,3-dihydro-1H-pyrrolizine-1-carboxylic acid 2. Biological Function and Applications: a COX1 and COX2 inhibitor, potent analgesic activity, a non-steroidal anti-inflammatory drug mainly used for its potent analgesic properties in the short-term management of post-operative pain, and in eye drops to relieve ocular itching associated with seasonal allergic conjunctivitis 3. Chemical Compound Classifications: a cyclooxygenase 2 inhibitor, a cyclooxygenase 1 inhibitor, a non-steroidal anti-inflammatory drug and an analgesic 4. Conjugate Base (if applicable): a ketorolac(1-) 5. Conjugate Acid (if applicable): N/A 6. Origin (if applicable): N/A
B.8 Task 8: Biology - Protein Description Generation
B.8.1 Task Definition
Protein description generation, also known as protein captioning, refers to creating an accurate, informative, and useful description of a protein given its amino acid sequence. These sequences are composed of letters representing individual amino acids and are the standard linear representations of protein primary structure. The objective is to generate a natural language description that captures essential biological characteristics of the protein, such as its function, cellular location, family or domain classification, and any relevant structural or catalytic properties. An example of a sequence-description pair is shown in Tab. 30. In using LLMs to complete this task, the input is an amino acid sequence (e.g., MKWVTFISLLFLFSSAYSRGVFRRDTH…) and the output is a protein description [71].
B.8.2 Task Significance
Protein sequences are long, complex, and typically require expert analysis and database lookups (e.g., UniProt [^22] https://www.uniprot.org, PDB [^23] https://www.rcsb.org/) to determine their biological functions. Automating the generation of descriptions helps make protein information more accessible and supports tasks like genome annotation, database curation, and bioinformatics research. Recent tools such as AnnoPRO [72] apply deep learning to predict protein function from sequences, reducing the need for manual curation and accelerating biological discovery [72].
B.8.3 Data Acquisition and Preprocessing
We collected protein sequences and their corresponding descriptions from the SciKnowEval dataset, which sources its protein and caption entries from the UniProtKB database [^24] https://www.uniprot.org/help/uniprotkb —a comprehensive, curated resource for protein sequence and functional information [73]. Reference descriptions are written by experts, curated by UniProt, and serve as the human reference for this task. From this dataset, we filtered a subset of 100 high-quality samples for our task. These samples were selected on length, all of which are over 900 characters.
B.8.4 Illustrative Example
An illustrative example is shown in Tab. 30 All samples for T8 are publicly available.
The sample input is an amino acid string. The human reference captures key biological attributes, including functional roles (e.g., antimicrobial activity), cellular localization (extracellular), molecular interactions (lipopolysaccharide binding), and structural features (e.g., $\beta$-strands with supporting PDB annotations). This example demonstrates the level of specificity and scientific grounding expected in generated protein descriptions.
Considering contextual grounding, we adopt a role-based prompting strategy. The model is instructed to assume the role of a protein researcher tasked with writing descriptions for protein sequences and encouraged to use domain-specific terminology with the model prompt shown in Tab. 31.
Table 30: T8BioPDG- A sample of an protein in its amino acid sequence and its corresponding description.
Amino acid sequence (Sample input): “METQRASLCLGRWSLWLLLLGLVVPSASAQALSYREAVLRAVDRLNEQSSEANLYRL LELDQPPKADEDPGTPKPVSFTVKETVCPRPTRQPPELCDFKENGRVKQCVGTVTLD QIKDPLDITCNEVQGVRGGRLCYCRPRFCVCVGRG" Protein description (Human-written reference): This protein exhibits microbicidal activity and plays a crucial role in the antimicrobial humoral immune response mediated by antimicrobial peptides, defense against both Gram-negative and Gram-positive bacteria, and in the innate immune response. It is found in the extracellular space and has the ability to bind to lipopolysaccharides. It has a signal peptide that spans from amino acid 1 to 29, indicating it is directed outside the cell. Structurally, it contains beta-strands with evidence from PDB:2NC7, specifically in the regions from amino acids 135 to 139 and 142 to 146. This combination of functional, localization, and structural attributes, notably its role in immune responses, extracellular location, lipopolysaccharide-binding capacity, presence of a signal peptide, and beta-strand formation, makes it distinct within the protein universe.
Table 31: T8BioPDG- Model prompt.
You are a protein researcher in charge of writing descriptions of proteins given their sequence. Use domain specific terminology and specific molecule names. Be as specific as possible. Please provide your description in paragraph format. Here is the sequence:
B.8.5 Evaluation Rubric
We collaborated with two graduate students from the University of Michigan studying biology to create the evaluation rubric. We first presented an initial draft of the rubric, created using online sources, and fine-tuned it with the experts’ assistance across three separate meetings until both experts agreed with the rubric. The experts’ also provided us with sources that explained their reasoning for including or removing certain items in the rubric. The rubric comprises five checklist items that captures the key information present in a protein description. The detailed checklist items are listed as follows:
  1. 1.
    Domains/Motifs: Functional regions (e.g., kinase domain, zinc finger) [74, 75].
  • Functional characteristics
    1. 2.
      Functional Role: How a molecule contributes to biological systems, including the processes it is involved in (e.g., metabolism, cellular functions) and the specific role it plays (e.g., enzymatic, structural, or signaling) [76, 77].
    2. 3.
      Cellular Localization: "Cellular location (e.g., mitochondrial matrix, cell membrane).
    3. 4.
      Gene Ontology: Identify key GO terms.
  1. 5.
    Interactions: The physical and functional associations within a cell or organism, including but not limited to protein partners, ligands/substrates, and cofactors [76, 77].
B.8.6 Checklist-mapped Reference
Consistent with the approach in T1, we generated checklist-mapped references for both human and model outputs. Because T8 has only five checklist items, a single model output extraction prompt was sufficient. The item-level extraction prompt in Tab. 32 and the model output extraction prompt are available in our public GitHub repository. The checklist-mapped reference for the example in Tab. 30 is shown in Tab. 33.
Table 32: T8BioPDG- Prompt for extracting checklist-mapped reference.
You are assisting a protein researcher in extracting key information from a protein description. Given a description, extract its Primary Structure: Amino acid sequence (e.g., 141 residues in $\alpha$-globin). Extract crucial related information as completely as possible. Extractions should come directly from the description in full sentence(s). If no related information is mentioned in the description, state "N/A" (as a string). This is the protein description:
Table 33: T8BioPDG- Checklist-mapped reference.
1. Domains/Motifs: N/A 2. Functional Role: This protein exhibits microbicidal activity and plays a crucial role in the antimicrobial humoral immune response mediated by antimicrobial peptides, defense against both Gram-negative and Gram-positive bacteria, and in the innate immune response. 3. Cellular Localization: It is found in the extracellular space 4. Gene Ontology: N/A 5. Interactions: and has the ability to bind to lipopolysaccharides.
B.9 T9MedicalDR
B.9.1 Task Definition
The objective of T9 is to assess a model’s ability to infer a Primary Discharge Diagnosis (PDD) from parts of a SOAP note. The input consists of selected sections of a SOAP note: the Chief Complaint, History of Present Illness (HPI), Past Medical History, Family History, Physical Exam, and Pertinent Results. Any explicit mentions of the diagnosis are manually removed. Given this input, the model must output the correct diagnosis and provide reasoning based on textual evidence.
B.9.2 Task Significance
This task addresses a challenge in clinical decision-making: deriving accurate diagnoses from patient information. Errors in diagnosis can lead to misdiagnosis, delayed treatment, or inappropriate care, with potentially severe consequences for patient outcomes. By enabling LLMs models to support diagnostic inference in a structured and interpretable way, this task contributes to improving medical safety and clinical decision support systems.
B.9.3 Data Acquisition and Preprocessing
Table 34: T9MedicalDR - Distribution of diagnoses.
Diagnosis # Cases
Atrial Fibrillation 6
Adrenal Insufficiency 7
Hypertension 6
Alzheimer 6
Pneumonia 6
Stroke 6
Gastro-oesophageal Reflux Disease 5
Epilepsy 6
Hyperlipidemia 2
Asthma 6
Heart Failure 6
Peptic Ulcer Disease 6
Diabetes 6
Pulmonary Embolism 6
Migraine 4
Thyroid Disease 5
Tuberculosis 5
Aortic Dissection 6
The DiReCT dataset consists of 511 de-identified discharge summaries [78]. From each note, specific sections relevant to the diagnostic process were extracted: chief complaint, history of present illness, past medical history, family history, physical exam, and pertinent results. To ensure high quality data, reasoning annotations were performed by nine licensed clinical physicians and verified for accuracy and completeness by three senior medical experts. Each diagnosis is one of 25 disease categories across five high-level clinical domains.
DiReCT originally contained information that our task has filtered out. The human reference is structured as follows: {o: [z, r, d]}. o is the extracted observation from raw text, z is the rationale to explain why an observation can be related to a diagnosis d, r is the section (from one of input1-6) of the clinical note where o is extracted, and finally, d is the name of the diagnosis. To calculate the #Rubric items of the human reference in Tab. 1 of the main paper, we concatenated this dict into a string and found its token length.
We then selected a smaller, high-quality, and representative subset from the dataset with the criteria of diversity and difficulty described in §3.2. To maintain diversity, we aimed to sample equally from all 25 final diagnosis. This promotes a balanced distribution that supports generalizable evaluation across multiple clinical domains. We first selected six samples per diagnosis type, resulting in an initial pool of 150 samples. Next, we prioritized cases with longer and more complex reasoning chains, as these are more likely to challenge model capabilities in clinical inference. From the 150 samples, we identified 100 with the longest human-generated reference outputs, reflecting the depth of clinical reasoning involved. Finally, we manually verified that each sample met both diversity and difficulty thresholds. Tab. 34 shows the diagnosis distribution of our final 100 samples across 18 diagnoses.
B.9.4 Illustrative Example
Due to the proprietary nature of the data, we cannot provide a specific example. However, Tab. 35 displays the model prompt. The model is instructed to diagnose the patient and provide supporting evidence in a structured dict_reasoning format. To ensure alignment with human references and enable verification, we constrained the model to select from a predefined list of possible diagnoses. This was necessary because, during early testing, the model often generated overly specific diagnoses— "Hypertension in the setting of atrial fibrillation" instead of "Hypertension"—which could not be matched against the simpler human reference.
Table 35: T9MedicalDR - Model prompt.
You are an medical expert. You will review a clinical ”Note“ with 6 inputs and generate an output to diagnose the disease that the patient has. All possible disease options are in a list structure: [’Hypertension’, ’Tuberculosis’, ’Alzheimer’, ’Gastritis’, ’Stroke’, ’Peptic Ulcer Disease’, ’Pituitary Disease’, ’Multiple Sclerosis’, ’Adrenal Insufficiency’, ’Migraine’, ’Cardiomyopathy’, ’Asthma’, ’Upper Gastrointestinal Bleeding’, ’Diabetes’, ’Aortic Dissection’, ’Hyperlipidemia’, ’Epilepsy’, ’Atrial Fibrillation’, ’Gastro-oesophageal Reflux Disease’, ’Acute Coronary Syndrome’, ’Pneumonia’, ’Pulmonary Embolism’, ’COPD’, ’Thyroid Disease’, ’Heart Failure’] You will also output your reasoning behind the diagnosis in a dict of dicts structure called dict_reasoning {{o: [z,r,d]…}}. Key: (string) Observation (o) - The EXACT extracted observation from raw text/input. Value: (list of strings) z = The rationale to explain why the observation is related to the diagnosis (string) r = "inputX" where X is the input integer (1-6) of the clinical note where o is extracted. (string) d = name of the diagnosis. (string) Note that if you can’t find any "Observations" your output should be: . Your response will have the structure: "Diagnosis: " diagnosis dict_reasoning Here is the note:
B.9.5 Evaluation Rubric
In T9, we did not consult domain experts as in previous tasks, but created our rubric independently. For this task, we introduce the concept of global versus instance-level checklist items.
  • Global checklist items are evaluated for every sample, regardless of the specific content of the human reference. For example, the item Diagnosis is global—it is always present in the human reference and model output and therefore always evaluated.
  • Instance-level checklist items are conditional: they are only evaluated when the relevant content appears in the human reference. For example, if the reference mentions the observation “elevated heart rate,” this becomes the key for an {Evidence: Reasoning} checklist item. The corresponding Reasoning value might be “An elevated heart rate is a common symptom of an infection.” These items are included only when such content is explicitly present in the human reference.
Thus, the number of checklist items in the final rubric varies based on the length and detail of the human reference, but comprises of three general items:
  1. 1.
    Diagnosis (global)
  2. 2.
    Each Evidence (instance-level): An observation that is textually present in the input
  3. 3.
    Each Reasoning (instance-level): The explanation in why the Evidence supports to the Diagnosis.
B.9.6 Checklist-mapped Reference
Due to the proprietary nature of the data, we cannot provide a specific example for the checklist-mapped reference. A prompt for generating the checklist-mapped reference was also unnecessary, as it was produced directly from the model output.
B.10 T10FinanceESG
B.10.1 Task Definition
ESG reports detail a company’s environmental, social, and governance practices and performance. These documents are typically long and complex, addressing a broad spectrum of key ESG issues such as carbon emissions, labor practices, and board diversity. The objective of ESG report summarization is to distill this information into concise summaries that emphasize the most pertinent aspects of a company’s ESG performance for investors and stakeholders. For each company, the model is given both an ESG report from a third-party rating agency and the company’s self-published ESG report as input for generating the summary.
B.10.2 Task Significance
Environmental, social, and governance considerations are increasingly recognized as critical components of long-term corporate performance. In response to global sustainability initiatives—such as the United Nations Sustainable Development Goals [^25] https://unglobalcompact.org/ and the Paris Agreement [^26] https://www.un.org/en/climatechange/paris-agreement—governments, institutions, and consumers have placed growing emphasis on corporate ESG practices. These factors now influence consumer behavior, regulatory compliance, and capital allocation. As a result, investors are integrating ESG performance into portfolio design and risk assessment.
Given this context, ESG disclosures have become critical tools for evaluating a company’s long-term value and social responsibility. However, ESG reports are often comprehensive and cover different aspects of ESG, not all of which are equally important for every company. Automated summarization that prioritizes the most relevant key issues enables more efficient consumption of ESG information. This task supports more informed decision-making by helping stakeholders focus on the ESG factors most relevant to a company’s financial, operational, and reputational outcomes.
B.10.3 Data Acquisition and Preprocessing
We utilize ESG reports published by MSCI [^27] https://www.msci.com, a leading ESG rating agency. MSCI’s ESG research covers over 10,000 companies across a wide range of sectors and regions. For each company, MSCI provides a comprehensive ESG report that outlines the company’s performance on ESG key issues, which vary by sector. [^28] https://www.msci.com/our-solutions/esg-investing/esg-industry-materiality-map In addition to performance data, the reports include textual summaries of the most relevant ESG issues and provide an overall ESG rating.
To ensure a diverse sample, we select 20 sectors with the highest number of companies rated by MSCI and choose 5 companies from each sector. These companies are selected to represent a range of ESG ratings (AAA, AA, A, BBB, BB, B). For each selected company, we collect both the most recent MSCI ESG report and the self-published ESG report available on the company’s website. Both types of reports are typically in PDF format and are converted to text using PyMuPDF4LLM.
From the MSCI reports, we extract only the sections that describe the company’s performance on key issues, discarding content such as summaries. For company-issued reports, which may include additional content such as financial data, we extract only the pages that contain ESG-related content.
Reference summaries are obtained directly from MSCI’s web interface rather than the PDF reports to ensure accurate extraction.
B.10.4 Illustrative Example
Due to the proprietary nature of the data, we cannot provide a specific example. However, in Tab. 36, we provide the model prompt used in obtaining model outputs.
Table 36: T10FinanceESG - Model prompt.
You are an ESG analyst. You will read the ESG report and sustainability report of a company. Based on the reports, write a summary of the company’s ESG performance. The summary should focus on the company’s performance on key ESG issues, and should include key details (e.g., strategies and initiatives). Word limit: 300 words.
B.10.5 Evaluation Rubric
The ESG report summaries discuss the company’s performance on ESG key issues. While the ESG report of a company details its performance on ESG key issues that are important to the corresponding sector, a well-written summary should cover the most significant key issues relevant to the company and give a precise overview of the company’s performance on these issues. We build the evaluation rubrics based on MSCI’s ESG scoring framework, which includes 33 key issues across 10 themes. [^29] https://www.msci.com/documents/1296102/34424357/MSCI+ESG+Ratings+Methodology.pdf The discussion of key issues is sector-specific, [^30] https://www.msci.com/our-solutions/esg-investing/esg-industry-materiality-map and therefore in this task we have different rubrics for different data samples, as shown in Tab. 37.
Table 37: Selected sectors and their corresponding key issues.
Sector Key Issues
Banks Access to Finance; Consumer Financial Protection; Corporate Behavior; Corporate Governance; Financing Environmental Impact; Human Capital Development; Privacy & Data Security
Biotechnology Access to Health Care; Corporate Behavior; Corporate Governance; Human Capital Development; Product Safety & Quality; Toxic Emissions & Waste
Construction & Engineering Corporate Behavior; Corporate Governance; Health & Safety; Human Capital Development; Opportunities in Clean Tech
Diversified Financials Access to Finance; Carbon Emissions; Corporate Behavior; Corporate Governance; Human Capital Development
Electronic Equipment, Instruments & Components Chemical Safety; Controversial Sourcing; Corporate Behavior; Corporate Governance; Labor Management; Opportunities in Clean Tech
Food Products Corporate Behavior; Corporate Governance; Opportunities in Nutrition & Health; Packaging Material & Waste; Product Carbon Footprint; Product Safety & Quality; Raw Material Sourcing; Water Stress
Health Care Equipment & Supplies Carbon Emissions; Corporate Behavior; Corporate Governance; Human Capital Development; Product Safety & Quality
Health Care Providers & Services Carbon Emissions; Corporate Behavior; Corporate Governance; Labor Management; Privacy & Data Security; Product Safety & Quality
Industrial Machinery Corporate Behavior; Corporate Governance; Labor Management; Opportunities in Clean Tech; Toxic Emissions & Waste
Media & Entertainment Carbon Emissions; Corporate Behavior; Corporate Governance; Labor Management; Privacy & Data Security
Metals and Mining - Non-Precious Metals Biodiversity & Land Use; Carbon Emissions; Community Relations; Corporate Behavior; Corporate Governance; Health & Safety; Labor Management; Toxic Emissions & Waste; Water Stress
Pharmaceuticals Access to Health Care; Corporate Behavior; Corporate Governance; Human Capital Development; Product Safety & Quality; Toxic Emissions & Waste
Real Estate Development & Diversified Activities Corporate Behavior; Corporate Governance; Health & Safety; Opportunities in Green Building; Product Safety & Quality
Real Estate Management & Services Corporate Behavior; Corporate Governance; Human Capital Development; Opportunities in Green Building
Retail - Consumer Discretionary Chemical Safety; Corporate Behavior; Corporate Governance; Labor Management; Privacy & Data Security; Product Carbon Footprint; Raw Material Sourcing
Semiconductors & Semiconductor Equipment Controversial Sourcing; Corporate Behavior; Corporate Governance; Human Capital Development; Opportunities in Clean Tech; Water Stress
Software & Services Carbon Emissions; Corporate Behavior; Corporate Governance; Human Capital Development; Opportunities in Clean Tech; Privacy & Data Security
Specialty Chemicals Carbon Emissions; Chemical Safety; Corporate Behavior; Corporate Governance; Opportunities in Clean Tech; Toxic Emissions & Waste; Water Stress
Telecommunication Services Carbon Emissions; Corporate Behavior; Corporate Governance; Labor Management; Privacy & Data Security
Utilities Carbon Emissions; Corporate Behavior; Corporate Governance; Human Capital Development; Opportunities in Renewable Energy; Toxic Emissions & Waste
The company’s performance on each key issue is broken down into several underlying indicators. For example, one of the indicators for environment-related issues is the extent to which the company has established programs and initiatives to address the issue. We examine the ESG report summaries for each sector and observe that the granularity of the discussion of key issues usually shares the same level of detail as the underlying indicators. Thus, we take these indicators as the items to be included in the rubrics, and we obtain the descriptions of these indicators based on the raw data MSCI uses to assess the company’s performance on these indicators.
Due to the proprietary nature of the data, in the paper, we cannot provide the list of rubrics items and their descriptions.
B.10.6 Checklist-mapped Reference
For each sample, to map the human-written summary into the reference checklist, we prompt the checklist mapper with the instruction that contains all the rubric items and their descriptions for the sector the company belongs to. The items are not separately extracted, as it is easier for the checklist mapper to distinguish between items that originate from the same key issue when they are presented together. We show the prompt for checklist mapping in LABEL:tab:T11-checklist-mappedReference-prompt.
Due to the proprietary nature of the data, we cannot provide a specific example for the checklist-mapped reference.
major_key_issue=12
[…]
$\begin$
}[!htbp]
[…]
$\begin$